### v2

# Analytical exciton energies in monolayer transition-metal dichalcogenides

Abstract
We derive an analytical expression for $s$-state exciton energies in monolayer transition-metal dichalcogenides (TMDCs): $E_{\text{ns}}=-{\text{Ry}}^{*}\times P_{n}/{(n-1/2+0.479\,r^{*}_{0}/\kappa)^{2}}$, $n=1,2,...$, where $r^{*}_{0}$ and $\kappa$ are the dimensionless screening length and dielectric constant of the surrounding medium; $\text{Ry}^{*}$ is an effective Rydberg energy scaled by the dielectric constant and exciton reduced mass; $P_{n}(r^{*}_{0}/\kappa)$ is a function of variables $n$ and $r^{*}_{0}/\kappa$. Its values are around 1.0 so we consider it a term that corrects the Rydberg energy. Despite the simple form, the suggested formula gives exciton energies with high precision compared to the exact numerical solutions that accurately describe recent experimental data for a large class of TMDC materials, including WSe2, WS2, MoSe2, MoS2, and MoTe2. To achieve these results, we have developed a so-called regulated perturbation theory by combining the conventional perturbation method with several elements of the Feranchuk-Komarov operator method, including the Levi-Civita transformation, the algebraic calculation technique via the annihilation and creation operators, and the introduction of a free parameter to optimize the convergence rate of the perturbation series. This universal form of exciton energies could be helpful in various physical analyses, including retrieval of the material parameters such as reduced exciton mass and screening length from the available measured exciton energies.
Exciton, transition-metal dichalcogenides, analytical energy, regulated perturbation theory, FK operator method
Introduction – In the last decade, exciton energies in different monolayer materials have been measured and studied theoretically simultaneously [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. Pioneer studies [1, 2] indicate that unlike 3D bulk semiconductors, electrons and holes in 2D atomically thin films such as monolayer transition-metal dichalcogenides (TMDCs) couple to each other by Rytova-Keldysh interaction [17, 18, 19, 20] when forming neutral excitons. As this interaction behaves as logarithm potential $\ln(r)$ in the short-range regime and becomes a conventional Coulomb interaction $1/r$ in the long-range regime [18, 20], the exciton energies of low-lying states ($n<5$) do not obey the two-dimensional hydrogen atom’s Ryberg series. Thus, an accurate theory of exciton binding energies in 2D semiconductors is necessary to benchmark experiments.
Recently, different approaches have been developed to determine exciton binding energies when the reduced mass, screening length, and dielectric constant of the surrounding medium are known [12, 13, 14, 15]. In particular, our packages [13, 14, 15] based on the Feranchuk-Komarov operator method [21, 22] perform excellently in calculating exciton binding energies in monolayer TMDCs, even in a high magnetic field. On the other hand, some theoretical efforts were made to obtain analytical formulas for exciton binding energies in 2D semiconductors [5, 6, 7]. These formulas arise from modeling the electron-hole interaction potential in 2D semiconductors to make the associated Wannier Schrödinger equation analytically solvable. This tactic leads to easy-to-use formulas generalized from the Rydberg series with acceptable accuracy. However, of course, the accuracy of these formulas is limited by the accuracy of their modified electron-hole potentials. Therefore, a comprehensive theory is needed to describe exciton energies analytically.
In this Letter, we provide highly accurate analytical expressions for $s$-state exciton energies based on a novel approach, the regulated perturbation theory (RPT). The obtained formulas for exciton energies, simple but universal for a wide range of TMDC monolayers, are precise compared with the exact numerical solutions and describe the currently available experimental data well. This achievement is granted by the advances of the method. Unlike conventional perturbation theory, the RPT approach includes the Levi-Civita transformation, which allows for the algebraic calculation technique and introduces a free parameter to control the convergence rate of the perturbation correction series. Moreover, it does not need to modify the electron-hole potential and precisely calculate all matrix elements concerning the harmonic oscillator basis set.
Regulated perturbation theory – For energy spectra of an exciton in a TMDC monolayer, most studies consider the Schrödinger equation
$\left\{-\left(\frac{\partial^{2}}{\partial x^{2}}+\frac{\partial^{2}}{\partial y^{2}}\right)+{\hat{V}}_{h-e}(r)-E\right\}\psi(x,y)=0,$ (1)
[…]
where $\alpha=r_{0}/\kappa a_{0}^{*}$ is a dimensionless parameter dependent on the screening length $r_{0}$.
The work [15] suggests transforming the Schrödinger equation from $(x,y)$ to $(u,v)$ space by the Levi-Civita transformation [23]. The new equation in $(u,v)$ space,
$(\hat{H}-E\hat{R})\psi(u,v)=0,$ (3)
is excluded from the Coulomb singularity, making solving it more convenient. Notably, it allows the solving process to use the harmonic oscillator basis set $|n,m(\omega)\rangle$ with running indices $n=0,1,2,\cdots$ and $m=0,\pm 1,\pm 2,\cdots,\pm n$ (the principal and magnetic quantum numbers). The harmonic frequency $\omega$ can be used as a free parameter to regulate the convergence rate of the perturbation series so that the desired precision is granted. Moreover, the equation in $(u,v)$ space and the harmonic basis set can be presented via the creation and annihilation operators: $\hat{a}^{+},\hat{a}$ and $\hat{b}^{+},\hat{b}$. All matrix elements necessary for solving Eq. (3), such as $h_{k}=\langle n|\hat{H}|n\rangle$, $r_{k}=\langle n|\hat{R}|n\rangle$, $v_{jk}=\langle j|\hat{H}|k\rangle$, and $v^{\text{\tiny{\it R}}}_{jk}=\langle j|\hat{R}|k\rangle$ $(j\neq k)$, can then be calculated algebraically using the commutation relations $[\hat{a},\hat{a}^{+}]=1$ and $[\hat{b},\hat{b}^{+}]=1$. Here, we consider only $s$-states where $m=0$, so the basis set can be shortly written as $|n\rangle\equiv|n,m=0\rangle$. This algebraic approach obtained analytical forms of these matrix elements in Ref. [15]. We provide them in Appendix A for use in the present study.
Different from Ref. [15], where Eq. (3) was treated numerically with highly accurate exciton energies, we solve this equation analytically by the perturbation method, expanding wave functions and energies by the perturbation parameter $\beta$ as $\psi=\psi^{(0)}+\beta\Delta\psi^{(1)}+\beta^{2}\Delta\psi^{(2)}+\cdots$, $E=E^{(0)}+\beta\Delta E^{(1)}+\beta^{2}\Delta E^{(2)}+\cdots$. Plugging these expansions into Eq. (3), where the operators are separated as $\hat{H}=\hat{H}_{0}+\beta\,\hat{V}$ and $\hat{R}=\hat{R}_{0}+\beta\hat{V}_{R}$, we obtain formulas for the wave function $\psi^{(0)}$ and energy $E^{(0)}$ in the zero-order approximation and their corrections $\Delta\psi^{(s)}$, $\Delta E^{(s)}$ in any approximation order $s=1,2,3,\cdots$. The present study is interested in exciton energies only to the third-order approximation, whose formulas are
$E^{(0)}_{n}$ $=$ $\varepsilon_{n}=\frac{h_{n}}{r_{n}},\quad\quad\Delta E_{n}^{(1)}=0,$
$\Delta E_{n}^{(2)}$ $=$ $-\sum_{\begin{smallmatrix}j=0\\ j\neq n\end{smallmatrix}}^{+\infty}\dfrac{(v_{nj}-\varepsilon_{n}v^{\text{\tiny{\it R}}}_{nj})(v_{jn}-\varepsilon_{n}v^{\text{\tiny{\it R}}}_{jn})}{(h_{j}-\varepsilon_{n}r_{j}){r_{n}}},$ (4)
$\Delta E^{(3)}_{n}$ $=$ $\sum_{\begin{smallmatrix}k=0\\ k\neq n\end{smallmatrix}}^{+\infty}{\sum_{\begin{smallmatrix}j=0\\ j\neq k,n\end{smallmatrix}}^{+\infty}\dfrac{(v_{nk}-\varepsilon_{n}v^{\text{\tiny{\it R}}}_{nk})P_{kj}(v_{jn}-\varepsilon_{n}v^{\text{\tiny{\it R}}}_{jn})}{(h_{k}-\varepsilon_{n}r_{k})(h_{j}-\varepsilon_{n}r_{j}){r_{n}}}}.$
Here, we use the notation $P_{kj}=v_{kj}-\varepsilon_{n}v^{\text{\tiny{\it R}}}_{kj}$ for short; energies in the $s$-order approximation are calculated by equation $E^{(s)}=E^{(0)}+\Delta E^{(1)}+\Delta E^{(2)}+\cdots+\Delta E^{(s)}$.
Table 1 shows exciton energies in monolayer WSe2 in the zero-, second-, and third-order approximations for $1s$ to $5s$ states, calculated by formulas (Analytical exciton energies in monolayer transition-metal dichalcogenides) and compared with the numerically exact solutions [15]. The free parameter $\omega$ is chosen by the variational condition $\partial E^{(0)}/\partial\omega=0$ for these energies. As shown in the Table, the energy precision is very high, within 1.0 meV for the second-order approximation, while less than 0.2 meV for the third-order approximation. We note that the current limit of the experimental measurement is about 1.0 meV.
Table 1: Exciton energies (meV) in monolayer WSe2 encapsulated by hBN slabs with screening length $r_{0}=4.21$ nm, exciton reduced mass $\mu=0.19\,m_{e}$, and dielectric constant $\kappa=4.34$, calculated by Eqs. (Analytical exciton energies in monolayer transition-metal dichalcogenides) and compared to the exact numerical solutions [15].
$E^{(0)}$ $E^{(2)}$ $E^{(3)}$ $E_{num.}$ [15]
$1s$164.98 -168.09 -168.55 -168.60
$2s$39.46 -37.05 -38.52 -38.57
$3s$16.94 -16.22 -16.40 -16.56
$4s$9.32 -9.03 -9.06 -9.13
$5s$5.87 -5.73 -5.74 -5.77
We have also verified the accuracy of formulas (Analytical exciton energies in monolayer transition-metal dichalcogenides) numerically for some other TMDC monolayers: WS2, MoSe2, and MoS2. It shows the same high precision, less than $0.5$ meV, in the third-order approximation, which is not provided here but is available in Supplementary [24].
Analytical exciton energies – As shown above, formulas (Analytical exciton energies in monolayer transition-metal dichalcogenides) for the corrections of up to the third-order approximation give extremely accurate exciton energies with precision within 0.5 meV compared to the numerically exact solutions. Even for the energies in the second-order approximation, the precision is about 1.0 meV, a limit for experimental measurement. Therefore, we consider them for analytical exciton energies. Indeed, by substituting the explicit forms of matrix elements $h_{k}$, $r_{k}$, $v_{jk}$, and $v^{\text{\tiny{\it R}}}_{jk}$ from Appendix A into Eqs. (Analytical exciton energies in monolayer transition-metal dichalcogenides), we get particular exciton energies in the zero-order approximation as
$E^{(0)}_{n}=\frac{\kappa^{2}}{4{r_{0}^{*}}^{2}}\frac{1}{\lambda^{2}}-\frac{2\kappa}{r_{0}^{*}}f_{n}(\lambda).$ (5)
Here, $r_{0}^{*}=r_{0}/a_{0}^{*}$ is defined as a dimensionless screening length; $\lambda$ is a variational parameter defined from the free parameter $\omega$ by $\lambda=\kappa/\omega r_{0}^{*}$; $f_{n}(\lambda)$ is a function dependent on $\lambda$ only, whose explicit form is given in Appendix A. For convenience, from now on, the principle quantum number $n$ is counted from 1 (instead of zero as in the previous calculations) to 5, corresponding to the states from $1s$ to $5s$.
Formula (5) depends on parameter $\lambda$, which we consider a variational parameter and determine by the equation
$\frac{\partial E^{(0)}_{n}}{\partial\lambda}=0.$ (6)
We must solve this equation to get $\lambda$ and plug it into Eq. (5) to get the needed analytical energies. We demonstrate the solving process for $ns$ states ($n\leq 5$) in Supplementary [24] and provide only the main idea and final results in the main text.
Because of the complexity of function $f_{n}(\lambda)$, analytically solving Eq. (6) is not trivial, requiring the Taylor expansion of this function. From another side, numerically solving this equation for specific cases of monolayers WSe2, WS2, MoSe2, and MoS2 encapsulated by hBN slabs with the material parameters given in Table 3, we get particular solutions for $1s$ state: $\lambda=0.803,0.863,0.582,$ and $0.750$, corresponding to each material. These results suggest we expand function $f_{1}(\lambda)$ around the average value $\lambda_{0}=0.75$. This expansion leads Eq. (6) to a polynomial equation concerning an unknown $\delta\lambda=\lambda-0.75$, making it easy to get approximate analytical solutions consistent with the numerical ones. Analogically, we get $\lambda$ for other $ns$ states by expanding functions $f_{n}(\lambda)$ around the points $\lambda_{0}=(n+1/2)/2$ and then solving Eq. (6). We combine all the solutions in one generalized formula as
$\lambda$ $=$ $\frac{1}{2}\left(n+0.5\right)+\delta\lambda,$ (7)
with
$\delta\lambda$ $=$ $\left(0.47\,{\kappa}/{r_{0}^{*}}-0.525\right)n\,.$ (8)
We note that the actual solutions of Eq. (6) are not precisely as given by formula (8); however, a slight modification has been made to get a general formula for all states $ns$. This tactic is reasonable because exact energies should not depend on the free parameter, and its modification around the extremum point does not change the solutions much. Reference [22] provides a detailed discussion of how to choose such a free parameter.
Plugging $\lambda$ from Eq. (7) into Eq. (5), we get exciton energies in analytical form, however, with more than 5% precision compared to the exact numerical solutions. So, for a higher precision, we will add the second-order correction of the perturbation theory to the energy (5), which does not change its form. After some analytical transformations and reasonable truncations of the polynomials, whose details are provided in Supplement [24], we finally obtain exciton energy for the $ns$ state as
$E_{\text{ns}}$ $=$ $-\frac{{\text{Ry}}^{*}\times P_{n}}{(n-0.5+0.479\,r_{0}^{*}/\kappa)^{2}}.$ (9)
Here, function $P_{n}$ has the form
$P_{n}(r_{0}^{*}/\kappa)=\frac{\exp{(a_{n}\,\delta)}}{1+b_{n}\,\delta}$ (10)
with $\delta=(0.47\,{\kappa}/{r_{0}^{*}}-0.525)$ and coefficients $a_{n}$ and $b_{n}$ defined as
$a_{n}$ $=$ $4.3/n-{6.6}/{n^{2}}+{3.2}/{n^{3}},$
$b_{n}$ $=$ $5.3/n-{7.6}/{n^{2}}+{4.0}/{n^{3}}\,.$ (11)
$P_{n}$ is a function of variable $r_{0}^{*}/\kappa$ only with values around 1.0 and $P_{n}\rightarrow 1$ when $n\rightarrow+\infty$, so we can interpret it as a term correcting the effective Rydberg energy Ry.
Table 2 gives exciton energies for the $ns$ states ($n\leq 5$) calculated by the analytical formula (9) compared with the exact numerical solutions for monolayer TMDCs [15]. We note that Ref. [15] provides only exciton energies for WSe2 and WS2; however, following the method given in this reference, we also calculate exact numerical exciton energies for MoSe2 and MoS2 to benchmark our analytical energies. The accuracy of the formula (9) is high, about 1.0 meV (an experimental precision limit) or less.
Table 2: Exciton energies of $ns$ states calculated by the analytical formula (9) compared with the numerical exact solutions in Ref. [15] for monolayer TMDCs encapsulated by hBN.
WSe2 WS2 MoSe2 MoS2
$r_{0}^{*}/\kappa$ 0.803 0.719 1.333 0.893
$E_{\text{1s}}$ (meV) -167.96 -177.80 -229.88 -219.34
Ref. [15] -168.60 -178.62 -231.96 -220.18
$E_{\text{2s}}$ (meV) -37.94 -39.12 -58.71 -50.83
Ref. [15] -38.57 -39.73 -60.63 -51.81
$E_{\text{3s}}$ (meV) -16.28 -16.61 -26.60 -22.04
Ref. [15] -16.56 -16.90 -27.31 -22.46
$E_{\text{4s}}$ (meV) -9.00 -9.13 -15.07 -12.25
Ref. [15] -9.13 -9.28 -15.41 -12.44
$E_{\text{5s}}$ (meV) -5.70 -5.77 -9.68 -7.78
Ref. [15] -5.77 -5.85 -9.87 -7.89
[…]
Validation with the available experimental data – To obtain analytical energies (9), we expand the equation around specific points $\lambda_{0}=(n+0.5)/2$ with the condition $\lambda\gg\delta\lambda$, which defines the working range of $r_{0}^{*}/\kappa$. From Eq. (8) for $\delta\lambda$, we get the following boundary for our theory
$0.89>r_{0}^{*}/\kappa\gg 0.46\quad\text{or}\quad 0.89<r_{0}^{*}/\kappa\ll 18.8$ (12)
Furthermore, we verify formula (9) by comparing it with experimental data for monolayers WSe2 [9], WS2, MoSe2, MoS2, and MoTe2 [10]. The comparison in Fig. S-1 demonstrates a high consistency between the experiments and theory. Experimental data of Liu et al. [8] for monolayer WSe2 also support our theory: $\Delta E_{2s-1s}=131$(130.3) meV, $\Delta E_{3s-2s}=21$ (21.7) meV, $\Delta E_{4s-2s}=30$ (29.0) meV, where the values in parenthesis are theoretical from Eq. (9).
Figure 1: Exciton energies for the $ns$ states in monolayers WSe2, WS2, MoSe2, MoS2, and MoTe2 calculated by formula (9) (horizontal axis) compared with the experimental data [10, 9] (vertical axis). Material parameters used for calculations are provided in Table 3.
Conclusion – A novel approach based on the perturbation theory has been developed, including some elements of the FK operator method, to achieve $s$-state exciton energies in monolayer transition-metal dichalcogenides. The third-order approximation gives almost exact solutions, while the second-order approximation of the perturbation theory gives exciton energies with high precision, within 1.0 meV compared to the exact numerical solutions. This accuracy suggests we use them for analytical exciton energies. Although simple and universal for $s$-states, the constructed analytical formula describes experimental exciton energies for a wide range of monolayer TMDCs. This general approach can be applied to other kinds of 2D semiconduction materials.
Acknowledgments
This work was funded by Vietnam Ministry of Education and Training under grant number B2022-SPS-09-VL and carried out by the high-performance cluster at Ho Chi Minh City University of Education, Vietnam.
A Material parameters and matrix elements
In this study, we use material parameters in Table 3
[…]
$\kappa$ $\mu$ ($m_{e}$) $r_{0}$ (nm) Ry (meV) $a^{*}_{0}$ (nm) Refs.
WSe2 4.34 0.19 4.21 137.24 1.209 [15]
WS2 4.16 0.175 3.76 137.59 1.258 [15]
MoSe2 4.40 0.35 3.90 245.97 0.665 [10]
MoS2 4.45 0.275 3.40 188.94 0.856 [10]
MoTe2 4.30 0.37 6.10 272.26 0.615 [24]
For calculating $s$-state exciton energies by perturbation theory formulas (2), we need the following matrix elements:
$h_{j}$ $=$ $\langle j|\hat{H}|j\rangle=\frac{\omega}{4}(2j+1)$
$\quad-\frac{2}{\omega\alpha}\left[(2j+1)u_{jj}+ju_{j,j-1}+(j+1)u_{j,j+1}\right],$
$r_{j}$ $=$ $\langle j|\hat{R}|j\rangle=\frac{1}{\omega}(2j+1),$ (13)
$v_{jk}$ $=$ $\langle j|\hat{H}|k\rangle\,\,(j\neq k)=\frac{\omega}{4}\left(k\delta_{j,k-1}+j\delta_{j,k+1}\right)$
$\quad-\frac{2}{\omega\alpha}\left[(2k+1)u_{j,k}+ku_{j,k-1}+(k+1)u_{j,k+1}\right],$
$v^{\text{\tiny R}}_{jk}$ $=$ $\langle j|\hat{R}|k\rangle=\frac{1}{\omega}\left(k\delta_{j,k-1}+j\delta_{j,k+1}\right),$ (14)
where
$u_{jk}(\lambda)$ $=$ $\sum_{s=0}^{\text{min}(k,j)}\sum_{t=0}^{j+k-2s}(-1)^{j+k+t}{{j+k-2s}\choose{t}}$ (15)
×($\qquad\qquad\times{{j}\choose{s}}{{k}\choose{s}}J_{2s+t+1}(\lambda).$
In Eq. (15)), we use the definite integrals
$J_{p}(\lambda)=\int\limits_{0}^{+\infty}\frac{dq}{(1+q)^{p}\sqrt{q^{2}+\lambda^{2}}}$ (16)
with $p\geq 1$ and $\lambda=1/\omega\alpha>0$. These integrals can be calculated numerically or by using the following iterative formulas
$J_{p}=\frac{(2p-3)J_{p-1}-(p-2)J_{p-2}+\lambda}{(\lambda^{2}+1)(p-1)}$ (17)
with $p\geq 2$, where $J_{1}(\lambda)$ has the explicit formula:
$J_{1}(\lambda)=\frac{\ln{\left(\lambda+\sqrt{\lambda^{2}+1}\right)}+\ln{\left(1+\sqrt{\lambda^{2}+1}\right)}-\ln(\lambda)}{\sqrt{\lambda^{2}+1}}.$ (18)
Noting that althought $J_{0}(\lambda)$ is disvergent, relation (17) is still valid for $p=2$ by considering the limit
${\textrm{lim}}_{p\rightarrow 0}\;pJ_{p}(\lambda)=1.$
Function $f_{n}(\lambda)$ in Eq. (3) is calculated by
$f_{n}(\lambda)=u_{n-1,n-1}+\frac{(n-1)u_{n-1,n-2}+nu_{n-1,n}}{2n-1}.$
Supplemental materials for:
Thermal effect on magnetoexciton energy spectra in monolayer
transition-metal dichalcogenides
Hanh T. Dinh, 1,† Ngoc-Hung Phan,1,† Duy-Nhat Ly,1, Dai-Nam Le,2, Ngoc-Tram D. Hoang,1 Nhat-Quang Nguyen,1 Phuoc-Thien Doan,1 and Van-Hoang Le 1
1Computational Physics Key Laboratory K002, Department of Physics, Ho Chi Minh City University of Education, Ho Chi Minh City 72759, Vietnam
2 Department of Physics, University of South Florida, Tampa, Florida 33620, USA
These authors contributed equally to this work.
Emails: hoanglv@hcmue.edu.vn (V.-H. Le).
S-I Algebraic form of the Schrödinger equation for an exciton in monolayer transition-metal dichalcogenides
S-I.1 The Schrödinger equation for an electron-hole system via the Levi-Civita transformation
The Schrödinger equation for the relative motion of electron and hole can be written in atomic units as
$\left\{-\left(\frac{\partial^{2}}{\partial x^{2}}+\frac{\partial^{2}}{\partial y^{2}}\right)+{\hat{V}}_{h-e}(r)-E\right\}\psi(x,y)=0,$ (S-1)
[…]
${\hat{V}}_{e-h}(x,y)=-2\int\limits_{0}^{+\infty}\frac{dq}{\sqrt{1+\alpha^{2}q^{2}}}\;\textrm{e}^{-qr},$ (S-2)
where $\alpha=r_{0}/\kappa a_{0}^{*}$. We note that most studies use the Keldysh potential expressed via Struve and Bessel functions, which are suitable only for numerical calculations.
The Levi-Civita transformation – For more effectiveness in solving proccess, we apply the Levi-Civita transformation
$x=u^{2}-v^{2},\;y=2uv,$ (S-3)
to the Schrödinger equation (S-1) to rewrite it in the $(u,v)$ space as
$\left\{-\frac{1}{4}\left(\frac{\partial^{2}}{\partial{u}^{2}}+\frac{\partial^{2}}{\partial{v}^{2}}\right)+\hat{V}_{K}(u,v)-E\left({u}^{2}+{v}^{2}\right)\right\}\psi(u,v)=0,$ (S-4)
where the interaction potential in the $(u,v)$ space is defined as ${\hat{V}}_{K}(u,v)=\left({u}^{2}+{v}^{2}\right){\hat{V}}_{h-e}$, which has the form
${\hat{V}}_{K}(u,v)=-2\int\limits_{0}^{+\infty}\frac{dq}{\sqrt{1+\alpha^{2}q^{2}}}\;\textrm{e}^{-q(u^{2}+v^{2})}(u^{2}+v^{2}).$ (S-5)
To obtain equations (S-4) and for further use, we also derive the distance $r$ and the angular momentum $\hat{l}_{z}$ in $(u,v)$ coordinates as follows
$r=u^{2}+v^{2},\quad{\hat{l}_{z}}=-\frac{i}{2}\left(v\frac{\partial}{\partial u}-u\frac{\partial}{\partial v}\right).$ (S-6)
For the application of the Levi-Civita transformation to two-dimensional atomic systems and calculation technique, one can also find more details in Ref. [23].
S-I.2 Algebraic formalism via annihilation and creation operators
We rewrite the Schrödinger equation (S-3) in the form
$\left({\hat{H}}-E{\hat{R}}\right)|\psi\rangle=0,$ (S-7)
where
${\hat{H}}=-\frac{1}{4}{\hat{T}}+{\hat{V}}_{K},\quad{\hat{T}}=\dfrac{\partial^{2}}{\partial{u}^{2}}+\dfrac{\partial^{2}}{\partial{v}^{2}},\quad{\hat{R}}={u}^{2}+{v}^{2}.$ (S-8)
Equation (S-6) is similar to those describing a two-dimensional anharmonic oscillator, suggesting we use its algebraic representation via annihilation and creation operators for convenience in calculations. We then use the harmonic osscilator’s wave functions to establish a basis set and calculate all necessary matrix elements by the algebraic method. More details of this algebraic approach can be found in monograph [22].
First, we define annihilation and creation operators as follows
$\hat{\alpha}=\sqrt{\frac{\omega}{2}}\left({u+\frac{1}{\omega}\frac{\partial}{{\partial u}}}\right),\quad\hat{\alpha}^{+}=\sqrt{\frac{\omega}{2}}\left({u-\frac{1}{\omega}\frac{\partial}{{\partial u}}}\right),$
$\quad\hat{\beta}=\sqrt{\frac{\omega}{2}}\left({v+\frac{1}{\omega}\frac{\partial}{{\partial v}}}\right),\quad\hat{\beta}^{+}=\sqrt{\frac{\omega}{2}}\left({v-\frac{1}{\omega}\frac{\partial}{{\partial v}}}\right).$ (S-9)
Here, $\omega$ in the above definition is usually an angular frequency of the harmonic oscillator. However, we can freely choose its value to manipulate the converge rate of the solutions, see [22] and references therein. Then, we use a canonical transform to get operators
$\hat{a}=\frac{1}{{\sqrt{2}}}\left({\hat{\alpha}-i\hat{\beta}}\right),\quad\hat{a}^{+}=\frac{1}{{\sqrt{2}}}\left({{{\hat{\alpha}}^{+}}+i{{\hat{\beta}}^{+}}}\right),\quad\hat{b}=\frac{1}{{\sqrt{2}}}\left({\hat{\alpha}+i\hat{\beta}}\right),\quad\hat{b}^{+}=\frac{1}{{\sqrt{2}}}\left({{{\hat{\alpha}}^{+}}-i{{\hat{\beta}}^{+}}}\right).$ (S-10)
Combinning two equations (S-I.2) and (S-10), the definition for the annihilation and creation operators is as follows
${\hat{a}}$ $=$ $\frac{\sqrt{\omega}}{2}\left(u-i\,v\right)+\frac{1}{2\,\sqrt{\omega}}\left(\frac{\partial}{\partial u}-i\,\frac{\partial}{\partial v}\right),$
${\hat{a}}^{+}$ $=$ $\frac{\sqrt{\omega}}{2}\left(u+i\,v\right)-\frac{1}{2\,\sqrt{\omega}}\left(\frac{\partial}{\partial u}+i\,\frac{\partial}{\partial v}\right),$
${\hat{b}}$ $=$ $\frac{\sqrt{\omega}}{2}\left(u+i\,v\right)+\frac{1}{2\,\sqrt{\omega}}\left(\frac{\partial}{\partial u}+i\,\frac{\partial}{\partial v}\right),$
${\hat{b}}^{+}$ $=$ $\frac{\sqrt{\omega}}{2}\left(u-i\,v\right)-\frac{1}{2\,\sqrt{\omega}}\left(\frac{\partial}{\partial u}-i\,\frac{\partial}{\partial v}\right).$ (S-11)
The purpose of definition (S-I.2) for operators $\hat{a}$, $\hat{a}^{+}$, $\hat{b}$, and $\hat{b}^{+}$ is twofold. First, these operators remain satisfying the commutation relations
$\left[{\hat{a}},{\hat{a}}^{+}\right]=1,\;\;\left[{\hat{b}},{\hat{b}}^{+}\right]=1,$ (S-12)
typical for annihilation and creation operators; and second, the angular momentum becomes a neutral operator as in the form
${\hat{l}_{z}}=\frac{1}{2}({\hat{a}}^{+}{\hat{a}}-{\hat{b}}^{+}{\hat{b}}).$ (S-13)
We also note that, operators ${\hat{a}}$, ${\hat{a}}^{+}$ are independent from operators ${\hat{b}}$, ${\hat{b}}^{+}$, meaning they commute with each other as
$\left[{\hat{a}},{\hat{b}}\right]=\left[{\hat{a}}^{+},{\hat{b}}^{+}\right]=\left[{\hat{a}},{\hat{b}}^{+}\right]=\left[{\hat{a}}^{+},{\hat{b}}\right]=0.$ (S-14)
Finally, using the annihilation and creation operators defined by equations (S-I.2), we can rewrite all the terms in the Schrödinger equation (S-7) as
${\hat{T}}=$ $\dfrac{\partial^{2}}{\partial{u}^{2}}+\dfrac{\partial^{2}}{\partial{v}^{2}}$ $=\;\omega\left({\hat{a}}{\hat{b}}+{{\hat{a}}}^{+}{{\hat{b}}}^{+}-{{\hat{a}}}^{+}{\hat{a}}-{{\hat{b}}}^{+}{\hat{b}}-1\right),$
${\hat{R}}=$ ${u}^{2}+{v}^{2}$ $=\;\frac{1}{\omega}\left({\hat{a}}{\hat{b}}+{{\hat{a}}}^{+}{{\hat{b}}}^{+}+{{\hat{a}}}^{+}{\hat{a}}+{{\hat{b}}}^{+}{\hat{b}}+1\right).$ (S-15)
The potential operator $\hat{V}_{K}$ can be expressed via $\hat{R}$ as
${\hat{V}}_{K}=-2\int\limits_{0}^{+\infty}\frac{dq}{\sqrt{1+\alpha^{2}q^{2}}}\;\textrm{e}^{-q\hat{R}}{\hat{R}}.$ (S-16)
S-I.3 Harmonic oscillator basic set and algebraically calculating matrix elements
Basis set – The Hamiltonian of two-dimentional isotropic harmonic oscillator associated with the annihilation and creation operators (S-I.2) has the algebraic form $\hat{H}_{osc}=\omega({\hat{a}}^{+}{\hat{a}}+{\hat{b}}^{+}{\hat{b}}+1)$. We will choose the eigenvectors of this Hamiltonian to be a basis set for our resolving process considering the Schrödinger equation (S-4) describing a two-dimensional anharmonic oscillator in the $(u,v)$- space. In addition, because the angular momentum conservation of excitons in monolayer TMDCs, we also require the basis set functions to be eigenfunctions of the operator ${\hat{l}}_{z}$ (Eq.(S-13)). As a results, we obtain
${|k,m\rangle}=\frac{1}{\sqrt{{(k+m)}!{(k-m)}!}}({\hat{a}}^{+})^{k+m}({\hat{b}}^{+})^{k-m}|0(\omega)\rangle$ (S-17)
as a basis set, where the vacuum state $|0(\omega)\rangle$ sastifies the equations:
${\hat{a}}\,{|0(\omega)\rangle}=0,\;\;{\hat{b}}\,|0(\omega)\rangle=0.$ (S-18)
Considering the algebraic form (S-13), we can verify that ${\hat{l}}_{z}\,{|k,m\rangle}=m\,{|k,m\rangle}$ with $m=0,\pm 1,\pm 2,...$. Therefore the running index $m$ of the basis set is fixed by the magnetic quantum number of the system. The other running index $k$ is integer with the values $k\geq|m|$. By using the commutation relations (S-12), (S-14) and equations (S-18), we can also certify the orthonormality of the basis set functions (S-17), i.e. $\langle k,m|k^{\prime},m^{\prime}\rangle=\delta_{k,k^{\prime}}\delta_{m,m^{\prime}},$ where $\delta_{ij}$ is the Kronecker delta.
Matrix elements – Commutation relations (S-12), (S-14) and equations (S-18) are useful for constructing algebraic calculation technique based on the annihilation and creation operators. This kind of algebraic techique is given in textbooks on quantum mechanics and monograph [22]. For all the operators in the Schrödinger equation (S-4), matrix elements with repect to the basis set (S-17) can be calculated and formulated analytically in Ref. [15]. In the present paper, we consider only $s$-states ($m=0$) so the basis set becomes
${|n\rangle}=\frac{1}{n!}({\hat{a}}^{+}{\hat{b}}^{+})^{n}|0(\omega)\rangle.$ (S-19)
(i) It is easy to derive the following equations
$({\hat{a}}^{+}{\hat{a}}+{\hat{b}}^{+}{\hat{b}}+1)\,{|n\rangle}=(2n+1)\,{|n\rangle},\quad{\hat{a}}^{+}{\hat{b}}^{+}{|n\rangle}={(n+1)}\,{|n+1\rangle},\quad{\hat{a}}{\hat{b}}\;{|n\rangle}={n}\;{|n-1\rangle},$ (S-20)
which lead to the matrix elements
$N_{jk}=\langle j|\,({\hat{a}}^{+}{\hat{a}}+{\hat{b}}^{+}{\hat{b}}+1)\,{|k\rangle}=(2k+1)\,\delta_{jk},$
$M_{jk}^{+}=\langle j|\,{\hat{a}}^{+}{\hat{b}}^{+}\,{|k\rangle}={(k+1)}\,\delta_{j,k+1},$
$M_{jk}=\langle j|\,{\hat{a}}{\hat{b}}\,{|k\rangle}={k}\,\delta_{j,k-1},$ (S-21)
useful for calculating matrix elements of the operators $\hat{T}$ and $\hat{R}$. We obtain
$\omega{R}_{jk}$ $=$ $\omega\,\langle j|\,{\hat{R}}\,{|k\rangle}=N_{jk}+M_{jk}+M^{+}_{jk}\qquad$ (S-22)
$=$ $(2k+1)\,\delta_{jk}+{k}\,\delta_{j,k-1}+{(k+1)}\,\delta_{j,k+1}\,,$
$\frac{1}{\omega}{T}_{jk}$ $=$ $\frac{1}{\omega}\,\langle j|\,{\hat{T}}\,{|k\rangle}=-N_{jk}+M_{jk}+M^{+}_{jk}$ (S-23)
$=$ $-(2k+1)\,\delta_{jk}+{k}\,\delta_{j,k-1}+{(k+1)}\,\delta_{j,k+1}\,.$
(ii) Differently, it is not trivial to calculate matrix elements of the operator $\hat{V}_{K}$. However, we can do it by using the technique of constructing operators in a normal form of annihilation and creation operators, given in Ref. [22]. Indeed, using the formulae (S-16), we have
$V^{\text{\tiny K}}_{jk}$ $=$ $\langle j|{\hat{V}_{K}}{|k\rangle}=(2k+1)\,U_{jk}+{k}\,U_{j,k-1}+{(k+1)}\,U_{j,k+1},$ (S-24)
where
$U_{jk}$ $=$ $-2\int\limits_{0}^{+\infty}\frac{dq}{\sqrt{1+{\omega}^{2}{\alpha}^{2}q^{2}}}\times\;\langle j|\,\textrm{e}^{-q\,(\hat{N}+\hat{M}+\hat{M}^{+})}|k\rangle.$ (S-25)
Here, we use the notations $\hat{N}={\hat{a}}^{+}{\hat{a}}+{\hat{b}}^{+}{\hat{b}}+1$, $\hat{M}={\hat{a}}{\hat{b}}$, and ${\hat{M}}^{+}={\hat{a}}^{+}{\hat{b}}^{+}$. To calculate matrix elements $U_{jk}$, we need the operator $\hat{O}=\textrm{e}^{-q\,(\hat{N}+\hat{M}+\hat{M}^{+})}$ in the normal form, in which the annihilation operator $\hat{M}$ in the right, the neutral operator $\hat{N}$ in the middle, and the creation operator ${\hat{M}}^{+}$ in the left. For operators creating a close algebra, there exists a procedure to create the normal form of the exponential operator [22]. In our case, we have the communication relations
$\left[\hat{M},\hat{N}\right]=2\hat{M},\quad\left[\hat{N},\hat{M}^{+}\right]=2\hat{M}^{+},\quad\left[\hat{M},\hat{M}^{+}\right]=\hat{N},$ (S-26)
that means operators $\hat{N}$, $\hat{M}$, and $\hat{M}^{+}$ creat a close algebra. Following this procedure given on page 232-233 of Ref. [22], we have the formula
$\hat{O}$ $=$ $\textrm{e}^{-q\,(\hat{N}+\hat{M}+\hat{M}^{+})}=e^{-\frac{q}{1+q}\,\hat{M}^{+}}e^{-\ln(1+q)\,\hat{N}}e^{-\frac{q}{1+q}\,\hat{M}}.$ (S-27)
Then, expanding the exponential operators in Eq.(S-27) by equation $e^{X}=\sum_{s=0}^{+\infty}\frac{1}{s!}X^{s}$ and using the fomulae obtained from Eqs.(S-20) as
$({\hat{M}}^{+})^{s}{|k\rangle}=\frac{(k+s)!}{k!}\,{|k+s\rangle},\quad{\hat{M}}^{s}\,{|k\rangle}=\frac{k!}{(k-s)!}\;{|k-s\rangle},\quad{\hat{N}}^{s}\,{|k\rangle}=(2k+1)^{s}\,{|k\rangle},$ (S-28)
we have
$O_{jk}=\langle j|\textrm{e}^{-q\,(\hat{N}+\hat{M}+\hat{M}^{+})}{|k\rangle}=(-1)^{j+k}\sum_{s=0}^{\text{min}(k,j)}{{j}\choose{s}}{{k}\choose{s}}\frac{q^{j+k-2s}}{(1+q)^{j+k+1}},$ (S-29)
where (${n\choose k}$ is the binomial coefficient, i.e. (${n\choose k}=\frac{n!}{(n-k)!k!}$.
Plugging Eq.(S-29) into (S-25), we obtain the final result for $U_{jk}$ as
$U_{jk}$ $=$ $-\frac{2}{\omega\,\alpha}\sum_{s=0}^{\text{min}(k,j)}\sum_{t=0}^{j+k-2s}(-1)^{j+k+t}{{j+k-2s}\choose{t}}{{j}\choose{s}}{{k}\choose{s}}$ (S-30)
×$\quad\qquad\qquad\qquad\qquad\quad\times\int\limits_{0}^{+\infty}\frac{dq}{(1+q)^{2s+t+1}\sqrt{q^{2}+1/\omega^{2}\alpha^{2}}}.$
In Eq. (S-30) the definite integral
$J_{p}(\lambda)=\int\limits_{0}^{+\infty}\frac{dq}{(1+q)^{p}\sqrt{q^{2}+\lambda^{2}}}$ (S-31)
with $p\geq 1$ and $\lambda=1/\omega\alpha>0$ is easy to calculate numerically. Besides, we can derive an iterative formula for this integral as follows:
$J_{p}=\frac{(2p-3)J_{p-1}-(p-2)J_{p-2}+\lambda}{(\lambda^{2}+1)(p-1)}$ (S-32)
with $p\geq 2$, where $J_{1}(\lambda)$ has the following explicit formula:
$J_{1}(\lambda)=\frac{\ln{\left(\lambda+\sqrt{\lambda^{2}+1}\right)}+\ln{\left(1+\sqrt{\lambda^{2}+1}\right)}-\ln(\lambda)}{\sqrt{\lambda^{2}+1}}.$ (S-33)
Noting that althought $J_{0}(\lambda)$ is disvergent, relation (S-32) is still valid for $p=2$ by considering the limit
${\textrm{lim}}_{p\rightarrow 0}\;pJ_{p}(\lambda)=1.$
so that
$J_{2}(\lambda)=\frac{J_{1}(\lambda)-1+\lambda}{\lambda^{2}+1}.$ (S-34)
For solving the Schrödinger equation (S-7) with the basis set (S-19), we need the following matrix elements:
$h_{j}=\langle j|\hat{H}|j\rangle,\quad v_{jk}=\langle j|\hat{H}|k\rangle\;(j\neq k),\quad r_{j}=\langle j|\hat{R}|j\rangle,\quad v^{\text{\tiny R}}_{jk}=\langle j|\hat{R}|k\rangle\;(j\neq k).$ (S-35)
From the above formulation, we can obtain these matrix elements explicitly as follows.
$h_{j}$ $=$ $\langle j|\hat{H}|j\rangle=-\frac{1}{4}T_{jj}+V^{\text{\tiny K}}_{jj}=\frac{\omega}{4}(2j+1)-\frac{2}{\omega\alpha}\left[(2j+1)u_{jj}+ju_{j,j-1}+(j+1)u_{j,j+1}\right],$
$v_{jk}$ $=$ $\langle j|\hat{H}|k\rangle=-\frac{1}{4}T_{jk}+V^{\text{\tiny K}}_{jk}\quad(j\neq k)$ (S-36)
$=$ $\frac{\omega}{4}\left(k\delta_{j,k-1}+j\delta_{j,k+1}\right)-\frac{2}{\omega\alpha}\left[(2k+1)u_{j,k}+ku_{j,k-1}+(k+1)u_{j,k+1}\right],$
$r_{j}$ $=$ $\langle j|\hat{R}|j\rangle=R_{jj}=\frac{1}{\omega}(2j+1),$
$v^{\text{\tiny R}}_{jk}$ $=$ $\langle j|\hat{R}|k\rangle=R_{jk}\quad(j\neq k)=\frac{1}{\omega}\left(k\delta_{j,k-1}+j\delta_{j,k+1}\right),$ (S-37)
where
$u_{jj}(\lambda)$ $=$ $\sum_{s=0}^{j}\sum_{t=0}^{2j-2s}(-1)^{t}{{2j-2s}\choose{t}}{{j}\choose{s}}^{2}J_{2s+t+1}(\lambda),$ (S-38)
$u_{jk}(\lambda)$ $=$ $\sum_{s=0}^{\text{min}(k,j)}\sum_{t=0}^{j+k-2s}(-1)^{j+k+t}{{j+k-2s}\choose{t}}{{j}\choose{s}}{{k}\choose{s}}J_{2s+t+1}(\lambda).$ (S-39)
S-II Regulated perturbation theory of solving the Schrödinger equation
S-II.1 Basic formulation of the perturbation theory
The Schrödinger equation considered in this study, Eq. (S-7), has the form
$\left({\hat{H}}-E{\hat{R}}\right)\psi(u,v)=0,$ (S-40)
where operators $\hat{H}$ and $\hat{R}$ are hermite. Noting that this equation was obtained from the Schrödinger equation in ($x,y$) space by the Levi-Civita transformation.
We solve equation (S-40) by the perturbation theory, deviding the operators as
$\hat{H}=\hat{H}_{0}+\beta\,\hat{V},\quad\hat{R}=\hat{R}_{0}+\beta\hat{V}_{R},$ (S-41)
with the perturbation parameter $\beta$. As shown in the following calculations, this parameter is introduced formally and does not effect the final results. The operators $\hat{H}_{0}$ and $\hat{R}_{0}$ are chosen so that they can be considered zero-order approximation of operators $\hat{H}$ and $\hat{R}$ and have common solutions $\psi^{(0)}$, $E^{(0)}$. Following the idea of the Feranchuk-Komarov (FK) operator method, we choose $\hat{H}_{0}$ and $\hat{R}_{0}$ so that their algebraic forms contain only neutral operators $\hat{a}^{+}\hat{a}$ and $\hat{b}^{+}\hat{b}$. It means these operators are diagonal respect to the basis set $|k\rangle$ ((S-18)). The remain parts, $\hat{V}$ and $\hat{V}_{R}$, are non-diagonal, containing $\hat{a}^{+}$, $\hat{a}$, $\hat{b}^{+}$, and $\hat{b}$ evenly. In this case, the matrix elements of theses operators with respect to the basis set (S-19) have the following properties
$\langle j|\hat{H}_{0}|j\rangle=\langle j|\hat{H}|j\rangle=h_{j},\quad\langle j|\hat{R}_{0}|j\rangle=\langle j|\hat{R}|j\rangle=r_{j}$
$\langle j|\hat{V}|k\rangle=\langle j|\hat{H}|k\rangle=v_{jk},\quad\langle j|\hat{V}_{R}|k\rangle=\langle j|\hat{R}|k\rangle=v^{\text{\tiny R}}_{jk},\quad(j\neq k),$ (S-42)
whose explicit expressions are given in Eqs.(S-I.3) and (S-I.3).
Wave function $\psi$ and energy $E$ are found in the expansions:
$\psi=\psi^{(0)}+\beta\Delta\psi^{(1)}+\beta^{2}\Delta\psi^{(2)}+\cdots=\psi^{(0)}+\sum_{s=1}^{+\infty}\beta^{s}\Delta\psi^{(s)}$
$E=E^{(0)}+\beta\Delta E^{(1)}+\beta^{2}\Delta E^{(2)}+\cdots=E^{(0)}+\sum_{s=1}^{+\infty}\beta^{s}\Delta E^{(s)}\;.$ (S-43)
Here, $\Delta\psi^{(s)}$ and $\Delta E^{(s)}$ are corrections on the wave function and energy in the $s$-order approximation.
Plugging expansions (S-II.1) into the Schrödinger equation (S-40), we obtain
$\left(\hat{H}_{0}-E^{(0)}\hat{R}_{0}\right)\psi^{(0)}+\beta\left[\left(\hat{H}_{0}-E^{(0)}\hat{R}_{0}\right)\Delta\psi^{(1)}+\left(\hat{V}-E^{(0)}\hat{V}_{R}-\Delta E^{(1)}\hat{R}_{0}\right)\psi^{(0)}\right]$
$\quad+\beta^{2}\left[\left(\hat{H}_{0}-E^{(0)}\hat{R}_{0}\right)\Delta\psi^{(2)}+\left(\hat{V}-E^{(0)}\hat{V}_{R}-\Delta E^{(1)}\hat{R}_{0}\right)\Delta\psi^{(1)}-\left(\Delta E^{(1)}\hat{V}_{R}+\Delta E^{(2)}\hat{R}_{0}\right)\psi^{(0)}\right]$
$\qquad+\sum_{s=3}^{+\infty}\beta^{s}\left[\left(\hat{H}_{0}-E^{(0)}\hat{R}_{0}\right)\Delta\psi^{(s)}+\left(\hat{V}-E^{(0)}\hat{V}_{R}-\Delta E^{(1)}\hat{R}_{0}\right)\Delta\psi^{(s-1)}\right.$
$\quad\quad\qquad\left.-\sum_{t=1}^{s-2}\left(\Delta E^{(s-t)}\hat{R}_{0}+\Delta E^{(s-t-1)}\hat{V}_{R}\right)\Delta\psi^{(t)}-\left(\Delta E^{(s)}\hat{R}_{0}+\Delta E^{(s-1)}\hat{V}_{R}\right)\psi^{(0)}\right]=0.$ (S-44)
From Eq. (S-II.1), we can get the equation for the zero-order approximation as
$\left(\hat{H}_{0}-E^{(0)}\hat{R}_{0}\right)\psi^{(0)}=0.$ (S-45)
Because of the diagonality of the operators $\hat{H}_{0}(\hat{a}^{+}\hat{a},\hat{b}^{+}\hat{b})$ and $\hat{R}_{0}(\hat{a}^{+}\hat{a},\hat{b}^{+}\hat{b})$, the eigen vectors and eigen values of Eq. (S-45) are as follows
$\psi^{(0)}_{n}=|n\rangle,\quad E^{(0)}_{n}=\frac{h_{n}}{r_{n}}.$ (S-46)
We can also get the equation for corrections in second-order approximation as
$\left(\hat{H}_{0}-E_{n}^{(0)}\hat{R}_{0}\right)\Delta\psi^{(1)}_{n}+\left(\hat{V}-E_{n}^{(0)}\hat{V}_{R}-\Delta E^{(1)}_{n}\hat{R}_{0}\right)|n\rangle=0.$ (S-47)
Multiflying the conjugate vector $|n\rangle$ and integrating over the space, we get
$\langle n|\left(\hat{H}_{0}-E_{n}^{(0)}\hat{R}_{0}\right)|\Delta\psi^{(1)}_{n}\rangle+\langle n|\left(\hat{V}-E^{(0)}_{n}\hat{V}_{R}-\Delta E^{(1)}_{n}\hat{R}_{0}\right)|n\rangle=0.$ (S-48)
Because of the hermite property, we rewrite the first term in Eq. (S-48) as $\langle\Delta\psi^{(1)}_{n}|(\hat{H}_{0}-E^{(0)}_{n})|n\rangle$ which means this first term vanishes. In other hand, from properties of matrix elements Eqs. (S-II.1), we have $\langle n|\hat{V}|n\rangle=0$ and $\langle n|\hat{V}_{R}|n\rangle=0$. As a consequence, Eq. (S-48) leads to $r_{n}\Delta E^{(1)}_{n}=0$. i.e.,
$\Delta E^{(1)}_{n}=0\,.$ (S-49)
Furthermore, Eq. (S-47) can be used to determine the correction to the wave function $\Delta\psi^{(1)}_{n}$. First, dealing with the second term in the equation, we have $\hat{V}|n\rangle=\sum_{j=0,j\neq n}^{+\infty}v_{jn}|j\rangle$ , $\hat{V}_{R}|n\rangle=\sum_{j=0,j\neq n}^{+\infty}v^{\text{\tiny R}}_{jn}|j\rangle$. Plugging these equation into Eq. (S-47), we obtain the following
$\Delta\psi^{(1)}_{n}=-\sum_{\begin{smallmatrix}j=0\\ j\neq n\end{smallmatrix}}^{+\infty}\frac{v_{jn}-E^{(0)}_{n}v^{\text{\tiny R}}_{jn}}{h_{j}-E^{(0)}_{n}r_{j}}\,|j\rangle=-\sum_{\begin{smallmatrix}j=0\\ j\neq n\end{smallmatrix}}^{+\infty}\frac{r_{n}v_{jn}-h_{n}v^{\text{\tiny R}}_{jn}}{r_{n}h_{j}-h_{n}r_{j}}\,|j\rangle\,.$ (S-50)
By this wav, we can get corrections to energies and wave functions for the second-order approximation as follows
$\Delta E^{(2)}_{n}=-\sum_{\begin{smallmatrix}j=0\\ j\neq n\end{smallmatrix}}^{+\infty}\frac{\left(v_{nj}-E^{(0)}_{n}v^{\text{\tiny R}}_{nj}\right)\left(v_{jn}-E^{(0)}_{n}v^{\text{\tiny R}}_{jn}\right)}{\left(h_{j}-E^{(0)}_{n}r_{j}\right){r_{n}}}=-\sum_{\begin{smallmatrix}j=0\\ j\neq n\end{smallmatrix}}^{+\infty}\frac{\left(r_{n}v_{nj}-h_{n}v^{\text{\tiny R}}_{nj}\right)\left(r_{n}v_{jn}-h_{n}v^{\text{\tiny R}}_{jn}\right)}{\left(r_{n}h_{j}-h_{n}r_{j}\right){r_{n}}^{2}}\,,$ (S-51)
$\Delta\psi^{(2)}_{n}$ $=$ $\sum_{\begin{smallmatrix}k=0\\ k\neq n\end{smallmatrix}}^{+\infty}\sum_{\begin{smallmatrix}j=0\\ j\neq k\\ j\neq n\end{smallmatrix}}^{+\infty}\frac{\left(v_{jn}-E^{(0)}_{n}v^{\text{\tiny R}}_{jn}\right)\left(v_{kj}-E^{(0)}_{n}v^{\text{\tiny R}}_{kj}\right)}{\left(h_{j}-E^{(0)}_{n}r_{j}\right)\left(h_{k}-E^{(0)}_{n}r_{k}\right)}\,|k\rangle$ (S-52)
$=$ $\sum_{\begin{smallmatrix}k=0\\ k\neq n\end{smallmatrix}}^{+\infty}\sum_{\begin{smallmatrix}j=0\\ j\neq k\\ j\neq n\end{smallmatrix}}^{+\infty}\frac{\left(r_{n}v_{jn}-h_{n}v^{\text{\tiny R}}_{jn}\right)\left(r_{n}v_{kj}-h_{n}v^{\text{\tiny R}}_{kj}\right)}{\left(r_{n}h_{j}-h_{n}r_{j}\right)\left(r_{n}h_{k}-h_{n}r_{k}\right)}\,|k\rangle.$
Considering Eq. (S-II.1), we can obtain corrections to energies and wave functions in any order of approximation. However, we are interested only corrections to energies in the approximation of up to the third order as
$\Delta E^{(3)}_{n}$ $=$ $\sum_{\begin{smallmatrix}k=0\\ k\neq n\end{smallmatrix}}^{+\infty}\sum_{\begin{smallmatrix}j=0\\ j\neq k\\ j\neq n\end{smallmatrix}}^{+\infty}\frac{\left(v_{nk}-E^{(0)}_{n}v^{\text{\tiny R}}_{nk}\right)\left(v_{kj}-E^{(0)}_{n}v^{\text{\tiny R}}_{kj}\right)\left(v_{jn}-E^{(0)}_{n}v^{\text{\tiny R}}_{jn}\right)}{\left(h_{j}-E^{(0)}_{n}r_{j}\right)\left(h_{k}-E^{(0)}_{n}r_{k}\right){r_{n}}}$ (S-53)
$=$ $\sum_{\begin{smallmatrix}k=0\\ k\neq n\end{smallmatrix}}^{+\infty}\sum_{\begin{smallmatrix}j=0\\ j\neq k\\ j\neq n\end{smallmatrix}}^{+\infty}\frac{\left(r_{n}v_{nk}-h_{n}v^{\text{\tiny R}}_{nk}\right)\left(r_{n}v_{kj}-h_{n}v^{\text{\tiny R}}_{kj}\right)\left(r_{n}v_{jn}-h_{n}v^{\text{\tiny R}}_{jn}\right)}{\left(r_{n}h_{j}-h_{n}r_{j}\right)\left(r_{n}h_{k}-h_{n}r_{k}\right){r_{n}}^{2}}\,.$
S-II.2 Numerical results
The analytical formulae above by the perturbation theory, Eqs. (S-46), can be used to obtain the s-state energies in the zero-order approximation. Moreover, the corrections to energies in the approximation of up to third order can be calculated by Eqs. (S-49), (S-51), and (S-53). We note that all the matrix elements in these formulae are given analytically by Eqs. (S-I.3) and (S-I.3). We will calculate energies for $1s$, $2s$, $3s$, $4s$, and $5s$ for monolayer TMDCs (WSe2, WS2, MoSe2, and MoS2) whose material parameters are given in Table S-I.
[…]
$\kappa$ $\mu$ ($m_{e}$) $r_{0}$ (nm) Ry (meV) $a^{*}_{0}$ (nm) Refs.
WSe2 4.34 0.19 4.21 137.24 1.211 [15]
WS2 4.16 0.175 3.76 137.59 1.256 [15]
MoSe2 4.40 0.35 3.90 245.97 0.664 [10]
MoS2 4.45 0.275 3.40 188.94 0.854 [10]
We have the s-state energies in zero-order approximation by Eq. (S-45) as
$E^{(0)}_{\text{ns}}=\frac{h_{n-1}}{r_{n-1}},$ (S-54)
where $n=1,2,3,4,5$ cosresponding to the states $1s,2s,3s,4s,5s$. Using the analytical formulae (S-I.3) and (S-I.3) for $h_{j}$ and $r_{j}$, we get
$E^{(0)}_{\text{ns}}=\frac{1}{4}\omega^{2}-\frac{2}{\alpha}f_{n}(\lambda)=\frac{1}{4\alpha^{2}\lambda^{2}}-\frac{2}{\alpha}f_{n}(\lambda).$ (S-55)
Here, instead of parameter $\omega$, we use the parameter $\lambda=1/\omega\alpha$; functions $f_{n}(\lambda)$ dependent only on variable $\lambda$ are defined by
$f_{n}(\lambda)=u_{n-1,n-1}+\frac{n-1}{2n-1}u_{n-1,n-2}+\frac{n}{2n-1}u_{n-1,n}.$ (S-56)
Plugging $\alpha=r_{0}/\kappa a_{0}^{*}$ into the formula, we have
$E^{(0)}_{\text{ns}}=\frac{\kappa^{2}}{4{r_{0}^{*}}^{2}\lambda^{2}}-\frac{2\kappa}{r_{0}^{*}}f_{n}(\lambda).$ (S-57)
where $r_{0}^{*}=r_{0}/a_{0}^{*}$ is a dimensionless screening length (in an unit of effective Bohr radius scaled by the dielectric constant). There is a free parameter $\lambda$ in energies (S-57) which we need to define. In principle, numerical exact solutions do not depend on this parameter, but approximate solutions are dependent on the choice of this parameter. Previous studies [22, 15] showed that choosing the free parameter around the solution of equation
$\frac{\partial E^{(0)}_{n}}{\partial\lambda}=0.$ (S-58)
is optimal, leading to accurate energies within some lowlying corrections. We choose the parameter $\lambda$ from Eq. (S-58) and put it to expression (S-57) to get energy in the zer0-order approximation. By numerical calculations with formulas (S-51) and (S-53), we obtain also exciton energies in second- and third-order approximations. Material parameters are picked up from Table S-I. Results in Tables S-II, S-III, S-IV, and S-V for monolayers WSe2, WS2, MoSe2, MoS2 encapsulated in hBN slabs demonstrate extremely high precision, within 1.0 meV in second-order approximation and 0.5 meV in third-order approximation.
Table S-II: Exciton energies (meV) calculated by regulated perturbation theory for monolayer WSe2 encapsulated by hBN slabs with screening length $r_{0}=4.21$ nm, exciton reduced mass $\mu=0.19\,m_{e}$, dielectric constant $\kappa=4.34$.
$E^{(0)}$ $E^{(2)}$ $E^{(3)}$ $E_{\text{num.}}$ [15]
$1s$164.98 -168.09 -168.55 -168.60
$2s$39.46 -37.05 -38.52 -38.57
$3s$16.94 -16.22 -16.40 -16.56
$4s$9.32 -9.03 -9.06 -9.13
$5s$5.87 -5.73 -5.74 -5.77
Table S-III: Exciton energies (meV) calculated by regulated perturbation theory for monolayer WS2 encapsulated by hBN slabs with screening length $r_{0}=3.76$ nm, exciton reduced mass $\mu=0.175\,m_{e}$, dielectric constant $\kappa=4.16$.
$E^{(0)}$ $E^{(2)}$ $E^{(3)}$ $E_{\text{num.}}$ [15]
$1s$174.66 -177.92 -178.40 -178.62
$2s$40.46 -38.25 -39.64 -39.73
$3s$17.21 -16.55 -16.74 -16.90
$4s$9.43 -9.16 -9.20 -9.28
$5s$5.93 -5.79 -5.80 -5.85
Table S-IV: Exciton energies (meV) calculated by regulated perturbation theory for monolayer MoSe2 encapsulated by hBN slabs with screening length $r_{0}=3.90$ nm, exciton reduced mass $\mu=0.35\,m_{e}$, dielectric constant $\kappa=4.40$.
$E^{(0)}$ $E^{(2)}$ $E^{(3)}$ $E_{\text{num.}}$ [15]
$1s$226.62 -231.01 -231.72 -231.96
$2s$62.73 -57.05 -60.20 -60.63
$3s$28.27 -26.53 -26.72 -27.31
$4s$15.88 -15.19 -15.14 -15.41
$5s$10.13 -9.79 -9.74 -9.87
Table S-V: Exciton energies (meV) calculated by regulated perturbation theory for monolayer MoS2 encapsulated by hBN slabs with screening length $r_{0}=3.40$ nm, exciton reduced mass $\mu=0.275\,m_{e}$, dielectric constant $\kappa=4.45$.
$E^{(0)}$ $E^{(2)}$ $E^{(3)}$ $E_{\text{num.}}$ [15]
$1s$215.29 -219.38 -220.00 -220.18
$2s$53.09 -49.52 -51.66 -51.81
$3s$23.01 -21.94 -22.18 -22.46
$4s$12.71 -12.28 -12.27 -12.44
$5s$8.03 -7.82 -7.82 -7.89
S-III Analytical expression for $s$-state exciton energies
Exciton energy for the $1s$ state
For 1s state, Eq. (S-57) become
$E^{(0)}_{1s}(r_{0}^{*}/\kappa)=\frac{\kappa^{2}}{4{r_{0}^{*}}^{2}}\frac{1}{\lambda^{2}}-\frac{2\kappa}{r_{0}^{*}}f_{1}(\lambda),$ (S-59)
where
$f_{1}(\lambda)=\frac{J(\lambda)+\lambda-1}{\lambda^{2}+1},\quad J(\lambda)=\frac{ln{\left[\left(1+\sqrt{\lambda^{-2}+1}\right)\left(1+\sqrt{\lambda^{2}+1}\right)\right]}}{\sqrt{\lambda^{2}+1}}.$ (S-60)
Here, $\lambda$ can be considered as a variational parameter and consequently is determined by the equation
$\frac{\partial E^{(0)}_{1s}}{\partial\lambda}=0.$ (S-61)
Basically, solving Eqs. (S-61), we can get an analytical formula for the $1s$-state energy. However, because of the complecity of function $J(\lambda)$ given by Eq. (S-60) the solving process is not trivial, requiring the Taylor expansion of this function. From another side, numerically solving Eq. (S-61) for the specific case of monolayers WSe2, WS2, MoSe2, and MoS2 encapsulated by hBN slabs with the material parameters given in Tab. S-I), we get $\lambda_{1}=0.803,0.863,0.582,$ and $0.750$, corresponding to each material. These results suggest us to expand function $f_{1}(\lambda)$ around the value $\lambda_{0}=0.75$. This point is obtained by the formula $\frac{1}{2}(n+1/2)$ with $n=1$. The expansion is
$f_{1}(\lambda)\simeq 2.454-6.399\,\lambda+11.956\,\lambda^{2}-14.594\,\lambda^{3}+11.079\,\lambda^{4}-4.757\,\lambda^{5}+0.884\,\lambda^{6}\quad.$ (S-62)
Polynomial (S-62) is truncated at $\lambda^{6}$ because its value for $\lambda$ in the range of $[0.5\rightarrow 0.9]$ is approximate to the function $f_{1}(\lambda)$ with an error less than $0.1\%$.
Plugging expansion (S-62) for $f_{1}(\lambda)$ into Eq. (S-61), we get an equation in the polynomial form for determining the parameter $\lambda$. The equation in term of $\delta\lambda=\lambda-0.75$ is given as follows
$-\frac{\kappa}{r^{*}_{0}}+1.120+2.326\,\delta\lambda+\cdots=0.$ (S-63)
Approximately, we consider Eq. (S-63) in the first order of $\delta\lambda$ which leads to the solution for $\delta\lambda$ as
$\delta\lambda$ $=$ $0.482\displaystyle 0.430\frac{\kappa}{r^{*}_{0}}-$ 0.482 . (S-64)
Substituting the material data in Tab. S-I into formula (S-64), we obtain the values of $\lambda=0.75+\delta\lambda$ as following: $0.804$ (0.803 WSe2), $0.865$ (0.863 WS2), $0.589$ (0.582 MoSe2), and $0.748$ (0.750 MoS2), which are consistent with numerically solving Eq. (S-61) in direct (the values given in the parenthesis).
Using the calculated above formulae, we rewrite the formula (S-59) for $1s$ state exciton energy in term of $\delta\lambda$ as
$E^{(0)}_{1s}$ $=$ $\frac{\kappa^{2}}{{r_{0}^{*}}^{2}}\frac{1}{(1.5+2\delta\lambda)^{2}}-\frac{2\kappa}{r_{0}^{*}}\left(0.757-0.664\delta\lambda\right.$ (S-65)
$\quad\left.+0.639\delta\lambda^{2}-0.654\delta\lambda^{3}+0.701\delta\lambda^{4}-0.778\delta\lambda^{5}+0.884\delta\lambda^{6}\right).$
From Eq. (S-64), we obtain a formula $\kappa/r^{*}_{0}=(\delta\lambda+0.482)/0.430$, which can be substituted into the second term of Eq. (S-65) to simplify this analytical exciton energy. As a result, formula (S-65) becomes
$E^{(0)}_{1s}$ $=$ $-\frac{\kappa^{2}}{{r_{0}^{*}}^{2}}\frac{2.042}{(1.5+2\,\delta\lambda)^{2}}\left(1-0.426\,\delta\lambda\right.$ (S-66)
$\quad\left.+1.30\,\delta\lambda^{2}-2.956\,\delta\lambda^{3}+6.318\,\delta\lambda^{4}-13.242\,\delta\lambda^{5}+27.63\,\delta\lambda^{6}\right).$
Numerical calculations of formula (S-66) give exciton energies with a precision of upto 5% compared with the exact numerical solutions. So we will add to this formula the correction of second order of perturbation theory. Especially, the correction of second-order approximation does not change the form of energy (S-66), but corrects only the coefficients in the equation. Indeed, putting the necessary calculated matrix elements into Eqs. (S-51) and then using the same Taylor expansion as above, we get
$E^{(2)}_{1s}$ $=$ $-\frac{\kappa^{2}}{{r_{0}^{*}}^{2}}\frac{2.08989}{(1.5+2\,\delta\lambda)^{2}}P_{1}(\delta\lambda),$ (S-67)
$P_{1}(\delta\lambda)$ $=$ $1-0.175\,\delta\lambda+0.742\,\delta\lambda^{2}-1.670\,\delta\lambda^{3}+3.363\,\delta\lambda^{4}-6.531\,\delta\lambda^{5}+12.579\,\delta\lambda^{6}.$
For the second-order approximation energy (S-67), we modify $\delta\lambda$ (Eq. (S-64)) slightly to get a general formula for all excited $s$-states as
$\delta\lambda_{n}$ $=$ $\left(0.470\frac{\kappa}{r^{*}_{0}}-0.525\right)n.$ (S-68)
This modification does not affect the final results much.
Polynomial $P_{1}(\delta\lambda)$ in Eq. (S-67) is approximately equal to one because of the smallness of the derivation $\delta\lambda$ ($0.062$ WSe2, $0.128$ WS2, $-0.173$ MoSe2, and $0.0003$ MoS2). To evaluate this polynomial, we show in Tab. S-VI its values for different monolayer TMDCs. We can see that, the polynomial has almost the value of 1.0 for MoS2 and can be truncated at $\delta\lambda^{0}$. For monolayers WSe2, this polynomial is equal to 0.99 and truncated at $\delta\lambda^{1}$. The same for monolayer WS2, the polynomial has the value of 0.987 but truncated at $\delta\lambda^{3}$. The worst result is for monolayer MoSe2 where the polynomial has the value of 1.066 and is truncated at $\delta\lambda^{5}$.
Table S-VI: Values of the polynomial $P_{1}(\delta\lambda)$ with the derivation $\delta\lambda$ taken from the materal data for WSe2, WS2, MoSe2, and MoS2. $P_{1}^{(N)}(\delta\lambda)$ means the polynomial $P_{1}(\delta\lambda)$ truncated at the term $\sim{\delta\lambda}^{N}$.
WSe2 WS2 MoSe2 MoS2
$\delta\lambda$ 0.062 0.128 -0.173 0.0003
$P_{1}^{(1)}(\delta\lambda)$ 0.989 0.9775 1.0303 0.99994
$P_{1}^{(2)}(\delta\lambda)$ 0.992 0.9897 1.0525 0.99994
$P_{1}^{(3)}(\delta\lambda)$ 0.9917 0.9862 1.0611 0.99994
$P_{1}^{(4)}(\delta\lambda)$ 0.9917 0.9871 1.0641 0.99994
$P_{1}^{(5)}(\delta\lambda)$ 0.9917 0.9869 1.0651 0.99994
$P_{1}^{(6)}(\delta\lambda)$ 0.9917 0.9869 1.0655 0.99994
Now, plugging the explicit form of $\delta\lambda$ Eq. (S-68) with $n=1$ into Eq. (S-67) and using the truncation discussed above, we can rewrite the analytical energy (S-67) in a general form for exciton energies as
$E^{(2)}_{1s}$ $=$ $-\frac{\text{Ry}^{*}\times P_{1}}{(0.650+0.311\,r^{*}_{0}/\kappa)^{2}},$ (S-69)
where $P_{1}$ is a polynomial whose values is around 1.0 as provided in Tab. S-VI. From the other hand, the energy spectra of two-dimensional hydrogen atom has the form
$E_{n}$ $=$ $-\frac{\text{Ry}}{(n-1/2)^{2}}.$ (S-70)
For the $1s$ state, $n=1$, so we expect the number $1-0.5=0.5$ instead of 0.569 in the denominator of Eq. (S-69). To make the formula (S-69) mimic the 2D hydrogen atom energy, we transform the main term in (S-69) as follows:
$\frac{2.08989}{(1.5+2\,\delta\lambda)^{2}}$ $=$ $\frac{2.08989}{(1.5+(2-x)\,\delta\lambda)^{2}}\times\frac{1}{\left(1+\dfrac{x\delta\lambda}{1.5+(2-x)\,\delta\lambda}\right)^{2}}.$
Thus, the exciton energy now become
$E^{(2)}_{1s}=\frac{\kappa^{2}}{{r_{0}^{*}}^{2}}\frac{2.08989\,{\text{Ry}^{*}}}{(1.5+(2-x)\,\delta\lambda)^{2}}\times P_{1}^{*}(\delta\lambda)$ (S-71)
with the polynomial $P_{1}$ modified as
$P_{1}^{*}(\delta\lambda)$ $=$ $\frac{P_{1}(\delta\lambda)}{\left(1+\dfrac{x\delta\lambda}{1.5+(2-x)\,\delta\lambda}\right)^{2}}.$ (S-72)
Chosing the value $x=0.462$ and putting the explicit form of $\delta\lambda$ (Eq. (S-68) into the denominator of (S-71), we finally obtain
$E^{(2)}_{1s}$ $=$ $-\frac{\text{Ry}^{*}\times P^{*}_{1}}{(0.5+0.479\,r^{*}_{0}/\kappa)^{2}}$ (S-73)
with the polynomial
$P^{*}_{1}(\delta\lambda)$ $=$ $1-0.791\,\delta\lambda+1.766\,\delta\lambda^{2}-3.636\,\delta\lambda^{3}+7.274\,\delta\lambda^{4}-14.323\,\delta\lambda^{5}+27.987\,\delta\lambda^{6}$
having values around 1.0 as shown in Table S-VII.
Table S-VII: Values of the polynomial $P^{*}_{1}(\delta\lambda)$ with the derivation $\delta\lambda$ taken from the materal data for WSe2, WS2, MoSe2, and MoS2. $P_{1}^{(N)}(\delta\lambda)$ means the polynomial $P^{*}_{1}(\delta\lambda)$ truncated at the term $\sim{\delta\lambda}^{N}$.
WSe2 WS2 MoSe2 MoS2
$\delta\lambda$ 0.062 0.128 -0.173 0.0003
$P_{1}^{(1)}(\delta\lambda)$ 0.9511 0.8986 1.1368 0.9997
$P_{1}^{(2)}(\delta\lambda)$ 0.9579 0.9276 1.1897 0.9997
$P_{1}^{(3)}(\delta\lambda)$ 0.9570 0.9200 1.2085 0.9997
$P_{1}^{(4)}(\delta\lambda)$ 0.9571 0.9219 1.2150 0.9997
$P_{1}^{(5)}(\delta\lambda)$ 0.9571 0.9214 1.2172 0.9997
$P_{1}^{(6)}(\delta\lambda)$ 0.9571 0.9216 1.2180 0.9997
From Table S-VII, we can see that polynomial $P^{*}_{1}$ truncated at $\sim\delta\lambda^{4}$ is accurate within a precision of 0.2%. Table S-VIII gives exciton energies for the $1s$ state calculated by analytical formula (S-73) compared with the exact numerical solutions. The accuracy of the formula is very hight, less than 1.0%.
Table S-VIII: Exciton energy of $1s$ state calculated by analytical formula (S-73). Comparison with the exact numerical solution [15] shows the high accuracy within errors less than 1%.
WSe2 WS2 MoSe2 MoS2
$r^{*}_{0}/\kappa$ 0.803 0.719 1.333 0.893
Ry (meV) 137.24 137.59 245.97 188.94
Formula(S-73) (meV) -168.03 -177.79 -230.79 -219.35
Exact num. [15] (meV) -168.60 -178.62 -231.96 -$220.18$
Relative error (%) 0.34 0.46 0.50 0.38
Analytical energies of $2s$ state
For 2s state, function $f_{2}(\lambda)$ has the form
$f_{2}(\lambda)$ $=$ $\frac{(1-3\lambda^{2}+\lambda^{4})J(\lambda)}{(\lambda^{2}+1)^{3}}+\frac{-6+11\lambda+8\lambda^{2}-3\lambda^{3}-\lambda^{4}+\lambda^{5}}{3(\lambda^{2}+1)^{3}}.$ (S-74)
By the same procedure as for $1s$-state, we obtain exciton energy for $2s$-state as
$E^{(2)}_{2s}$ $=$ $-\frac{\kappa^{2}}{{r_{0}^{*}}^{2}}\frac{1.325\,{\text{Ry}^{*}}}{(2.5+2\,\delta\lambda)^{2}}\times P_{2}(\delta\lambda),$ (S-75)
where
$P_{2}(\delta\lambda)$ $=$ $1-0.0398\,\delta\lambda+0.0953\,\delta\lambda^{2}-0.119\,\delta\lambda^{3}+0.131\,\delta\lambda^{4}-0.127\,\delta\lambda^{5}+0.137\,\delta\lambda^{6}.$ (S-76)
The equation $\partial E^{0}/\partial\lambda=0$ leads to $\delta\lambda=0.88\kappa/{r_{0}^{*}}-1.04$. However, we modify it slightly following Eq. (S-68) to get
$\delta\lambda$ $=$ $1.05\displaystyle 0.94\frac{\kappa}{r^{*}_{0}}-$ 1.05 . (S-77)
Put it into Eq. (S-75) we obtain
$E^{(2)}_{2s}$ $=$ $-\frac{\text{Ry}^{*}}{(1.633+0.348\,r^{*}_{0}/\kappa)^{2}}\times P_{2}(\delta\lambda),$ (S-78)
where values of $P_{2}$ given in Table S-IX are around 1.0.
Table S-IX: Values of the polynomial $P_{2}(\delta\lambda)$ with the derivation $\delta\lambda$ taken from the materal data for WSe2, WS2, MoSe2, and MoS2. $P_{2}^{(N)}(\delta\lambda)$ means the polynomial $P_{2}(\delta\lambda)$ truncated at the term $\sim{\delta\lambda}^{N}$.
WSe2 WS2 MoSe2 MoS2
$\delta\lambda$ 0.123 0.256 -0.346 0.0007
$P_{2}^{(1)}(\delta\lambda)$ 0.9951 0.9898 1.0138 0.99997
$P_{2}^{(2)}(\delta\lambda)$ 0.9965 0.9961 1.0252 0.99997
$P_{2}^{(3)}(\delta\lambda)$ 0.9963 0.9941 1.0301 0.99997
$P_{2}^{(4)}(\delta\lambda)$ 0.9963 0.9946 1.0320 0.99997
$P_{2}^{(5)}(\delta\lambda)$ 0.9963 0.9945 1.0326 0.99997
$P_{2}^{(6)}(\delta\lambda)$ 0.9963 0.9945 1.0328 0.99997
We use the following transformation
$\frac{1.325}{(2.5+2\,\delta\lambda)^{2}}$ $=$ $\frac{1.325}{(2.5+(2-x)\,\delta\lambda)^{2}}\times\frac{1}{\left(1+\dfrac{x\delta\lambda}{2.5+(2-x)\,\delta\lambda}\right)^{2}}$
to modify the form of Eq. (S-75) as
$E^{(2)}_{2s}=-\frac{\kappa^{2}}{{r_{0}^{*}}^{2}}\frac{1.325\,{\text{Ry}^{*}}}{(2.5+(2-x)\,\delta\lambda)^{2}}\times P_{2}^{*}(\delta\lambda)$ (S-79)
with the polynomial $P_{2}$ modified as
$P_{2}^{*}(\delta\lambda)$ $=$ $\frac{P_{2}(\delta\lambda)}{\left(1+\dfrac{x\delta\lambda}{2.5+(2-x)\,\delta\lambda}\right)^{2}}.$ (S-80)
Chosing the value $x=0.163$ and putting the explicit form of $\delta\lambda$ (Eq. (S-77) into the denominator of (S-79), we finally obtain
$E^{(2)}_{2s}$ $=$ $-\frac{\text{Ry}^{*}}{(1.5+0.496\,r^{*}_{0}/\kappa)^{2}}\times P^{*}_{2}$ (S-81)
with
$P^{*}_{2}(\delta\lambda)$ $=$ $1-0.171\,\delta\lambda+0.209\,\delta\lambda^{2}-0.226\,\delta\lambda^{3}+0.235\,\delta\lambda^{4}-0.231\,\delta\lambda^{5}+0.240\,\delta\lambda^{6}.$
Values of $P_{2}^{*}$ given in Table S-X are around 1.0.
Table S-X: Values of the polynomial $P^{*}_{2}(\delta\lambda)$ with the derivation $\delta\lambda$ taken from the materal data for WSe2, WS2, MoSe2, and MoS2. $P_{2}^{(N)}(\delta\lambda)$ means the polynomial $P_{2}(\delta\lambda)$ truncated at the term $\sim{\delta\lambda}^{N}$.
WSe2 WS2 MoSe2 MoS2
$\delta\lambda$ 0.124 0.256 -0.346 0.0007
$P_{2}^{(1)}(\delta\lambda)$ 0.9789 0.9563 1.0590 0.99989
$P_{2}^{(2)}(\delta\lambda)$ 0.9820 0.9701 1.0840 0.99989
$P_{2}^{(3)}(\delta\lambda)$ 0.9817 0.9663 1.0934 0.99989
$P_{2}^{(4)}(\delta\lambda)$ 0.9818 0.9673 1.0967 0.99989
$P_{2}^{(5)}(\delta\lambda)$ 0.9818 0.9670 1.0979 0.99989
$P_{2}^{(6)}(\delta\lambda)$ 0.9818 0.9671 1.0983 0.99989
From Table S-X, we can see that polynomial $P^{*}_{2}$ truncated at $\sim\delta\lambda^{3}$ is accurate within a precision of 0.2%. Table S-XI gives exciton energies for the $2s$ state calculated by analytical formula (S-81) compared with the exact numerical solutions.
Table S-XI: Exciton energy of $2s$ state calculated by analytical formula (S-81). Comparison with the exact numerical solution [15] shows the high accuracy.
WSe2 WS2 MoSe2 MoS2
$r^{*}_{0}/\kappa$ 0.803 0.719 1.333 0.893
Ry (meV) 137.24 137.59 245.97 188.94
Formula(S-81) (meV) -37.41 -38.59 -57.80 -50.03
Exact num. [15] (meV) -38.57 -39.73 -60.63 -51.81
Relative error (%) 3.0 2.9 4.7 3.4
Analytical energies of $3s$ state
For the $3s$ state, function $f_{3}(\lambda)$ has the form
$f_{3}(\lambda)$ $=$ $\frac{q_{3}(\lambda)J(\lambda)+p_{3}(\lambda)}{(\lambda^{2}+1)^{5}}$ (S-82)
with
$q_{3}(\lambda)$ $=$ $1-11\lambda^{2}+23.25\lambda^{4}-11\lambda^{6}+\lambda^{8},$
$p_{3}(\lambda)$ $=$ $-2.5+6.6\lambda+16.75\lambda^{2}-20.2\lambda^{3}-22.2\lambda^{4}+18.75\lambda^{5}+5.6\lambda^{6}-1.5\lambda^{7}-0.2\lambda^{8}+0.2\lambda^{9}.$
After the similar calculations, we get
$E^{(2)}_{3s}$ $=$ $-\frac{\kappa^{2}}{{r_{0}^{*}}^{2}}\frac{1.140\,\text{Ry}^{*}}{(3.5+2\,\delta\lambda)^{2}}\times P_{3}(\delta\lambda),$ (S-83)
with
$P_{3}(\delta\lambda)$ $=$ $1-0.0204\,\delta\lambda+0.0298\,\delta\lambda^{2}-0.0237\,\delta\lambda^{3}+0.111\,\delta\lambda^{4}-0.0478\,\delta\lambda^{5}+0.00772\,\delta\lambda^{6}.$
The equation $\partial E^{0}/\partial\lambda=0$ leads to $\delta\lambda=1.38\kappa/{r_{0}^{*}}-1.575$. However, we modify it slightly following Eq. (S-68) to get
$\delta\lambda$ $=$ $1.575\displaystyle 1.41\frac{\kappa}{r^{*}_{0}}-$ 1.575 (S-84)
Put it into Eq. (S-83), we obtain
$E^{(2)}_{3}$ $=$ $-\frac{\text{Ry}^{*}}{(2.642+0.327\,r^{*}_{0}/\kappa)^{2}}P_{3}(\delta\lambda),$ (S-85)
Values of $P_{3}(\delta\lambda)$ given in Table S-XII are around 1.0.
Table S-XII: Values of the polynomial $P_{3}(\delta\lambda)$ with the derivation $\delta\lambda$ taken from the materal data for WSe2, WS2, MoSe2, and MoS2. $P_{1}^{(N)}(\delta\lambda)$ means the polynomial $P_{2}(\delta\lambda)$ truncated at the term $\sim{\delta\lambda}^{N}$.
WSe2 WS2 MoSe2 MoS2
$\delta\lambda$ 0.185 0.364 -0.519 0.001
$P_{3}^{(1)}(\delta\lambda)$ 0.9962 0.9922 1.0106 0.99998
$P_{3}^{(2)}(\delta\lambda)$ 0.9972 0.9966 1.0186 0.99998
$P_{3}^{(3)}(\delta\lambda)$ 0.9971 0.9952 1.0219 0.99998
$P_{3}^{(4)}(\delta\lambda)$ 0.9972 0.9976 1.0299 0.99998
$P_{3}^{(5)}(\delta\lambda)$ 0.9972 0.9972 1.0317 0.99998
$P_{3}^{(6)}(\delta\lambda)$ 0.9972 0.9972 1.0319 0.99998
Further, we use the transformation
$\frac{1.140}{(3.5+2\,\delta\lambda)^{2}}$ $=$ $\frac{1.140}{(3.5+(2-x)\,\delta\lambda)^{2}}\times\frac{1}{\left(1+\dfrac{x\delta\lambda}{3.5+(2-x)\,\delta\lambda}\right)^{2}}.$
Thus, the exciton energy now become
$E^{(2)}_{3s}=\frac{\kappa^{2}}{{r_{0}^{*}}^{2}}\frac{1.140}{(3.5+(2-x)\,\delta\lambda)^{2}}\times P_{3}^{*}(\delta\lambda)$ (S-86)
with the polynomial $P_{3}$ modified as
$P_{3}^{*}(\delta\lambda)$ $=$ $\frac{P_{3}(\delta\lambda)}{\left(1+\dfrac{x\delta\lambda}{3.5+(2-x)\,\delta\lambda}\right)^{2}}.$ (S-87)
Chosing the value $x=0.107$ and putting the explicit form of $\delta\lambda$ (Eq. (S-84) into the denominator of (S-86), we finally obtain
$E^{(2)}_{3s}$ $=$ $-\frac{\text{Ry}^{*}}{(2.5+0.486\,r^{*}_{0}/\kappa)^{2}}\times P^{*}_{3},$ (S-88)
where
$P^{*}_{3}(\delta\lambda)$ $=$ $1-0.0816\,\delta\lambda+0.0670\,\delta\lambda^{2}-0.0473\,\delta\lambda^{3}+0.126\,\delta\lambda^{4}-0.0635\,\delta\lambda^{5}+0.0199\,\delta\lambda^{6}.$ (S-89)
Values given in Table S-XIII show that the function $P_{3}^{*}(\delta\lambda)$ is almost 1.0.
Table S-XIII: Values of the polynomial $P^{*}_{3}(\delta\lambda)$ with the derivation $\delta\lambda$ taken from the materal data for WSe2, WS2, MoSe2, and MoS2. $P_{3}^{(N)}(\delta\lambda)$ means the polynomial $P^{*}_{3}(\delta\lambda)$ truncated at the term $\sim{\delta\lambda}^{N}$.
WSe2 WS2 MoSe2 MoS2
$\delta\lambda$ 0.185 0.384 -0.519 0.001
${P_{3}^{*}}^{(1)}(\delta\lambda)$ 0.9849 0.9686 1.0424 0.99992
${P_{3}^{*}}^{(2)}(\delta\lambda)$ 0.9872 0.9785 1.0604 0.99992
${P_{3}^{*}}^{(3)}(\delta\lambda)$ 0.9869 0.9758 1.0670 0.99992
${P_{3}^{*}}^{(4)}(\delta\lambda)$ 0.9870 0.9786 1.0761 0.99992
${P_{3}^{*}}^{(5)}(\delta\lambda)$ 0.9870 0.9781 1.0785 0.99992
${P_{3}^{*}}^{(6)}(\delta\lambda)$ 0.9870 0.9781 1.0789 0.99992
Table S-XIV: Exciton energy of $3s$ state calculated by analytical formula (S-88) with polynomial $P^{*}_{3}$ by Eq. (S-89). Comparison with the exact numerical solution [15] shows the high accuracy.
WSe2 WS2 MoSe2 MoS2
$r^{*}_{0}/\kappa$ 0.803 0.719 1.333 0.893
Ry (meV) 137.24 137.59 245.97 188.94
Formula(S-88) (meV) -16.22 -16.58 -26.77 -21.94
Exact num. [15] (meV) -16.56 -16.90 -27.31 -22.46
Relative error (%) 2.1 1.9 2.0 2.3
Analytical energies of $4s$ state
For the $4s$ state, function $f_{4}(\lambda)$ has the form
$f_{4}(\lambda)$ $=$ $\frac{21q_{4}(\lambda)J(\lambda)+p_{4}(\lambda)}{21(\lambda^{2}+1)^{7}},$ (S-90)
where
$q_{4}(\lambda)$ $=$ $1-24\lambda^{2}+131.25\lambda^{4}-223.75\lambda^{6}+131.25\lambda^{8}-24\lambda^{10}+\lambda^{12}$
$p_{4}(\lambda)$ $=$ $-59.5+201\lambda+964.25\lambda^{2}-1696\lambda^{3}-3774.75\lambda^{4}+4625.75\lambda^{5}+4520.75\lambda^{6}$
$-3652.75\lambda^{7}-1759\lambda^{8}+1027.25\lambda^{9}+180\lambda^{10}-38.5\lambda^{11}-3\lambda^{12}+3\lambda^{13}.$
We obtain the exciton energy
$E^{(2)}_{4s}$ $=$ $-\frac{\kappa^{2}}{{r_{0}^{*}}^{2}}\frac{1.047}{(4.5+2\,\delta\lambda)^{2}}\times P_{4}(\delta\lambda),$ (S-91)
where
$P_{4}(\delta\lambda)$ $=$ $1-0.0106\,\delta\lambda+0.00858\,\delta\lambda^{2}-0.00749\,\delta\lambda^{3}+0.00017\,\delta\lambda^{4}-0.0118\,\delta\lambda^{5}+0.00086\,\delta\lambda^{6}.$
Solution of equation $\partial E^{0}/\partial\lambda=0$ leads to $\delta\lambda=1.88{\kappa}/{r^{*}_{0}}-2.09$; however, we slightly modify it for a general formulation to get
$\delta\lambda$ $=$ $2.1\displaystyle 1.88\frac{\kappa}{r^{*}_{0}}-$ 2.1 (S-92)
We ontain the exciton energy in the second-order approximation as
$E^{(2)}_{4s}$ $=$ $-\frac{1}{(3.66+0.305\,r^{*}_{0}/\kappa)^{2}}P_{4}(\delta\lambda),$ (S-93)
where the values of $P_{4}$ given in Table S-XV are around 1.0.
Table S-XV: Values of the polynomial $P_{4}(\delta\lambda)$ with the derivation $\delta\lambda$ taken from the materal data for WSe2, WS2, MoSe2, and MoS2. $P_{1}^{(N)}(\delta\lambda)$ means the polynomial $P_{2}(\delta\lambda)$ truncated at the term $\sim{\delta\lambda}^{N}$.
WSe2 WS2 MoSe2 MoS2
$\delta\lambda$ 0.247 0.512 -0.692 0.013
$P_{4}^{(1)}(\delta\lambda)$ 0.9974 0.9946 1.0074 0.9999
$P_{4}^{(2)}(\delta\lambda)$ 0.9979 0.9968 1.0115 0.9999
$P_{4}^{(3)}(\delta\lambda)$ 0.9978 0.9958 1.0139 0.9999
$P_{4}^{(4)}(\delta\lambda)$ 0.9978 0.9958 1.0140 0.9999
$P_{4}^{(5)}(\delta\lambda)$ 0.9978 0.9954 1.0159 0.9999
$P_{4}^{(6)}(\delta\lambda)$ 0.9978 0.9954 1.0159 0.9999
We make a transformation
$\frac{1.047}{(4.5+2\,\delta\lambda)^{2}}$ $=$ $\frac{1.047}{(4.5+(2-x)\,\delta\lambda)^{2}}\times\frac{1}{\left(1+\dfrac{x\delta\lambda}{4.5+(2-x)\,\delta\lambda}\right)^{2}}$
to get exciton energy
$E^{(2)}_{4s}=-\frac{\kappa^{2}}{{r_{0}^{*}}^{2}}\frac{1.047\,\text{Ry}^{*}}{(4.5+(2-x)\,\delta\lambda)^{2}}\times P_{4}^{*}(\delta\lambda)$ (S-94)
with the polynomial $P_{4}$ modified as
$P_{4}^{*}(\delta\lambda)$ $=$ $\frac{P_{4}(\delta\lambda)}{\left(1+\dfrac{x\delta\lambda}{4.5+(2-x)\,\delta\lambda}\right)^{2}}.$ (S-95)
Chosing the value $x=0.095$ and putting the explicit form of $\delta\lambda$ (Eq. (S-92) into the denominator of (S-94), we finally obtain
$E^{(2)}_{4s}$ $=$ $-\frac{\text{Ry}^{*}}{(3.5+0.488\,r^{*}_{0}/\kappa)^{2}}\times P^{*}_{4}$ (S-96)
with
$P^{*}_{4}(\delta\lambda)$ $=$ $1-0.0528\,\delta\lambda+0.0282\,\delta\lambda^{2}-0.0168\,\delta\lambda^{3}+0.0047\,\delta\lambda^{4}-0.0139\,\delta\lambda^{5}+0.0023\,\delta\lambda^{6}.$ (S-97)
Values of $P_{4}^{*}$ given in Table S-XVI are almost 1.0.
Table S-XVI: Values of the polynomial $P^{*}_{4}(\delta\lambda)$ with the derivation $\delta\lambda$ taken from the materal data for WSe2, WS2, MoSe2, and MoS2. $P_{1}^{(N)}(\delta\lambda)$ means the polynomial $P^{*}_{4}(\delta\lambda)$ truncated at the term $\sim{\delta\lambda}^{N}$.
WSe2 WS2 MoSe2 MoS2
$\delta\lambda$ 0.247 0.512 -0.692 0.0013
${P_{4}^{*}}^{(1)}(\delta\lambda)$ 0.9870 0.9729 1.0365 0.99993
${P_{4}^{*}}^{(2)}(\delta\lambda)$ 0.9887 0.9803 1.0501 0.99993
${P_{4}^{*}}^{(3)}(\delta\lambda)$ 0.9884 0.9781 1.0556 0.99993
${P_{4}^{*}}^{(4)}(\delta\lambda)$ 0.9884 0.9784 1.0566 0.99993
${P_{4}^{*}}^{(5)}(\delta\lambda)$ 0.9884 0.9779 1.0589 0.99993
${P_{4}^{*}}^{(6)}(\delta\lambda)$ 0.9884 0.9779 1.0591 0.99993
Table S-XVII: Exciton energy of $4s$ state calculated by analytical formula (S-96) with polynomial $P^{*}_{4}$ by Eq. (S-97). Comparison with the exact numerical solution [15] shows the high accuracy.
WSe2 WS2 MoSe2 MoS2
$r^{*}_{0}/\kappa$ 0.803 0.719 1.333 0.893
Ry (meV) 137.24 137.59 245.97 188.94
Formula(S-96) (meV) -8.96 -9.07 -15.12 -12.20
Exact num. [15] (meV) -9.13 -9.28 -15.41 -12.44
Relative error (%) 1.9 2.2 2.0 1.9
Analytical energies of $5s$ state
For the $5s$ state, we have
$f_{5}(\lambda)$ $=$ $\frac{9q_{5}(\lambda)J(\lambda)+p_{5}(\lambda)}{9(\lambda^{2}+1)^{9}},$ (S-98)
with
$q_{5}(\lambda)$ $=$ $1-42\lambda^{2}+436.75\lambda^{4}-1612.75\lambda^{6}+2462.27\lambda^{8}-1612.75\lambda^{10}+436.75\lambda^{12}-42\lambda^{14}+\lambda^{16}$
$p_{5}(\lambda)$ $=$ $-2.75+113\lambda+829.88\lambda^{2}-1854\lambda^{3}-6561.66\lambda^{4}+10369.25\lambda^{5}+18599.11\lambda^{6}-21400.75\lambda^{7}-21526.75\lambda^{8}$
$+18725.11\lambda^{9}+10285.25\lambda^{10}-6477.66\lambda^{11}-1890\lambda^{12}+865.88\lambda^{13}+104\lambda^{14}-18.75\lambda^{15}-\lambda^{16}+\lambda^{17}.$
Solution of equation $\partial E^{0}/\partial\lambda=0$ leads to $\delta\lambda=2.37\,{\kappa}/{r^{*}_{0}}-2.587$; however, we slightly modify it for a general formulation to get
$\delta\lambda$ $=$ $2.625\displaystyle 2.35\frac{\kappa}{r^{*}_{0}}-$ 2.625 . (S-99)
We have exciton energy for $5s$ state as
$E^{(2)}_{5s}$ $=$ $-\frac{\text{Ry}^{*}}{(4.723+0.251\,r^{*}_{0}/\kappa)^{2}}\times P_{5}(\delta\lambda),$ (S-100)
where
$P_{5}(\delta\lambda)$ $=$ $1-0.0034\,\delta\lambda+0.000005\,\delta\lambda^{2}-0.0046\,\delta\lambda^{3}+0.00027\,\delta\lambda^{4}-0.000006\,\delta\lambda^{5}+0.00007\,\delta\lambda^{6}.$ (S-101)
has the values around 1,0 shown in Table S-XVIII
Table S-XVIII: Values of the polynomial $P_{5}(\delta\lambda)$ with the derivation $\delta\lambda$ taken from the materal data for WSe2, WS2, MoSe2, and MoS2. $P_{5}^{(N)}(\delta\lambda)$ means the polynomial $P_{5}(\delta\lambda)$ truncated at the term $\sim{\delta\lambda}^{N}$.
WSe2 WS2 MoSe2 MoS2
$\delta\lambda$ 0.309 0.641 -0.865 0.0017
$P_{5}^{(1)}(\delta\lambda)$ 0.9990 0.9978 1.0029 0.99999
$P_{5}^{(2)}(\delta\lambda)$ 0.9990 0.9979 1.0029 0.99999
$P_{5}^{(3)}(\delta\lambda)$ 0.9988 0.9966 1.0059 0.99999
$P_{5}^{(4)}(\delta\lambda)$ 0.9988 0.9967 1.0060 0.99999
$P_{5}^{(5)}(\delta\lambda)$ 0.9988 0.9967 1.0060 0.99999
$P_{5}^{(6)}(\delta\lambda)$ 0.9988 0.9967 1.0060 0.99999
After the transformation
$\frac{0.9904}{(5.5+2\,\delta\lambda)^{2}}$ $=$ $\frac{0.9904}{(5.5+(2-x)\,\delta\lambda)^{2}}\times\frac{1}{\left(1+\dfrac{x\delta\lambda}{5.5+(2-x)\,\delta\lambda}\right)^{2}},$ (S-102)
the exciton energy becomes
$E^{(2)}_{5}=\frac{\kappa^{2}}{{r_{0}^{*}}^{2}}\frac{0.9904\,\text{Ry}^{*}}{(5.5+(2-x)\,\delta\lambda)^{2}}\times P_{5}^{*}(\delta\lambda)$ (S-103)
with the polynomial $P_{5}$ modified as
$P_{5}^{*}(\delta\lambda)$ $=$ $\frac{P_{5}(\delta\lambda)}{\left(1+\dfrac{x\delta\lambda}{5.5+(2-x)\,\delta\lambda}\right)^{2}}.$ (S-104)
Chosing the value $x=0.0944$ and putting the explicit form of $\delta\lambda$ (Eq. (S-99) into the denominator of (S-103), we finally obtain
$E^{(2)}_{5s}$ $=$ $-\frac{\text{Ry}^{*}}{(4.5+0.5001\,r^{*}_{0}/\kappa)^{2}}\times P^{*}_{5},$ (S-105)
where
$P^{*}_{5}(\delta\lambda)$ $=$ $1-0.0377\,\delta\lambda+0.0129\,\delta\lambda^{2}-0.0094\,\delta\lambda^{3}+0.0022\,\delta\lambda^{4}-0.0007\,\delta\lambda^{5}+0.00034\,\delta\lambda^{6}$ (S-106)
has values around 1.0 shown in Table S-XIX
Table S-XIX: Values of the polynomial $P^{*}_{5}(\delta\lambda)$ with the derivation $\delta\lambda$ taken from the materal data for WSe2, WS2, MoSe2, and MoS2. $P_{5}^{(N)}(\delta\lambda)$ means the polynomial $P^{*}_{5}(\delta\lambda)$ truncated at the term $\sim{\delta\lambda}^{N}$.
WSe2 WS2 MoSe2 MoS2
$\delta\lambda$ 0.309 0.641 -0.865 0.0017
${P_{5}^{*}}^{(1)}(\delta\lambda)$ 0.9884 0.9759 1.0326 0.99994
${P_{5}^{*}}^{(2)}(\delta\lambda)$ 0.9896 0.9812 1.0422 0.99994
${P_{5}^{*}}^{(3)}(\delta\lambda)$ 0.9893 0.9787 1.0483 0.99994
${P_{5}^{*}}^{(4)}(\delta\lambda)$ 0.9893 0.9791 1.0495 0.99994
${P_{5}^{*}}^{(5)}(\delta\lambda)$ 0.9893 0.9790 1.0485 0.99994
${P_{5}^{*}}^{(6)}(\delta\lambda)$ 0.9893 0.9790 1.0500 0.99994
Table S-XX: Exciton energy of $5s$ state calculated by analytical formula (S-105) with polynomial $P^{*}_{4}$ by Eq. (S-106). Comparison with the exact numerical solution [15] shows the high accuracy within errors less than 1%.
WSe2 WS2 MoSe2 MoS2
$r^{*}_{0}/\kappa$ 0.801 0.719 1.334 0.894
Ry (meV) 137.24 137.59 245.97 188.94
Formula(S-105) (meV) -5.65 -5.70 -9.67 -7.72
Exact num. [15] (meV) -5.77 -5.85 -9.87 -7.89
Relative error (%) 2.1 2.6 2.0 2.1
From Eqs. (S-73), (S-81), (S-88), (S-96), and (S-105) for analytical energies of $1s$, $2s$, $3s$, $4s$, and $5s$, we combine them into a general formula
$E_{ns}$ $=$ $-\frac{\text{Ry}^{*}}{(n-0.5+0.479\,r^{*}_{0}/\kappa)^{2}}\times P^{*}_{n}.$ (S-107)
Here, from different values in the denominators as 0.479, 0.496, 0.486, 0.488, 0.500 we choose one value 0.479 because this value makes the formula best fit with the exact numerical solutions. From different polynomials $P_{1}^{*}$, $P_{2}^{*}$, $P_{3}^{*}$, $P_{4}^{*}$, and $P_{5}^{*}$, we can unify them into one function as follows. First, we estimate the polynomials by functions as
$P^{*}_{1}(\delta\lambda)$ $=$ $1-0.791\,\delta\lambda+1.766\,\delta\lambda^{2}-3.636\,\delta\lambda^{3}+7.274\,\delta\lambda^{4}\sim\frac{\exp{(0.914\,\delta\lambda)}}{1+1.705\,\delta\lambda},$ (S-108)
$P^{*}_{2}(\delta\lambda)$ $=$ $1-0.171\,\delta\lambda+0.209\,\delta\lambda^{2}-0.226\,\delta\lambda^{3}+0.235\,\delta\lambda^{4}\sim\frac{\exp{(0.4545\,\delta\lambda)}}{1+0.625\,\delta\lambda},$ (S-109)
$P^{*}_{3}(\delta\lambda)$ $=$ $1-0.0816\,\delta\lambda+0.0670\,\delta\lambda^{2}-0.0473\,\delta\lambda^{3}+0.126\,\delta\lambda^{4}\sim\frac{\exp{(0.275\,\delta\lambda)}}{1+0.3566\,\delta\lambda},$ (S-110)
$P^{*}_{4}(\delta\lambda)$ $=$ $1-0.0528\,\delta\lambda+0.0282\,\delta\lambda^{2}-0.0168\,\delta\lambda^{3}\sim\frac{\exp{(0.179\,\delta\lambda)}}{1+0.2318\,\delta\lambda},$ (S-111)
$P^{*}_{5}(\delta\lambda)$ $=$ $1-0.0377\,\delta\lambda+0.0129\,\delta\lambda^{2}-0.0094\,\delta\lambda^{3}\sim\frac{\exp{(0.1183\,\delta\lambda)}}{1+0.156\,\delta\lambda}.$ (S-112)
Then, we rewrite them in one general function as
$P^{*}_{n}(r_{0}^{*}/\kappa)$ $=$ $\frac{\exp{(a_{n}\,\delta)}}{1+b_{n}\,\delta},$ (S-113)
where
$\delta=\frac{\delta\lambda}{n}=0.47\kappa/r_{0}^{*}-0.525,\quad a_{n}=\frac{4.3}{n}-\frac{6.6}{n^{2}}+\frac{3.2}{n^{3}},\quad b_{n}=\frac{5.3}{n}-\frac{7.6}{n^{2}}+\frac{4.0}{n^{3}}.$ (S-114)
Table S-XXI: Exciton energies of $ns$ states calculated by the analytical formula (S-115) compared with the numerical exact solutions in Ref. [15] for monolayer TMDCs encapsulated by hBN.
WSe2 WS2 MoSe2 MoS2
$r_{0}^{*}/\kappa$ 0.803 0.719 1.333 0.893
$E_{\text{1s}}$ (meV) -167.96 -177.80 -229.88 -219.34
Ref. [15] -168.60 -178.62 -231.96 -220.18
$E_{\text{2s}}$ (meV) -37.94 -39.12 -58.71 -50.83
Ref. [15] -38.57 -39.73 -60.63 -51.81
$E_{\text{3s}}$ (meV) -16.28 -16.61 -26.60 -22.04
Ref. [15] -16.56 -16.90 -27.31 -22.46
$E_{\text{4s}}$ (meV) -9.00 -9.13 -15.07 -12.25
Ref. [15] -9.13 -9.28 -15.41 -12.44
$E_{\text{5s}}$ (meV) -5.70 -5.77 -9.68 -7.78
Ref. [15] -5.77 -5.85 -9.87 -7.89
We can also compare analytical energies (S-115) with the formula suggested by Molas et al. [6] as
$E_{ns}^{Molas}$ $=$ $-\frac{\text{Ry}^{*}}{(n+\delta)^{2}}.$ (S-115)
Our formula is similar to the one of Molas with $\delta=-0.5+0.479*r_{0}^{*}/\kappa$ and $\text{Ry}^{*}(modify)={\text{Ry}}^{*}\times P_{n}$. For monolayer WSe2 in bHN, Molas’s formula is $E_{ns}^{Molas}=134{\text{meV}}/(n-0.099)^{2}$ while our formula is rewritten as $E_{ns}=134.5{\text{meV}}/(1-0.11)^{2}$. Notably, in our formula the effect of screening the charge is presented via the correction $P_{n}^{*}$.
S-IV Experimental validation
We use experimental data of Refs. [9, 10, 8] to test analytical formula (S-115). Some data are provided numericallly in the works, other data we extract from the figures of the works. Theoretically, we calculate exciton energies for monolayer Wse2, WS2, MoSe2, and MoS2 by formula (S-115) with material parameters picked up from Refs. [10] and [15]. For MoTe2, Ref. [10] give only two energy values: $E_{b}=177\pm 3$ meV and transition energy $2s-1s$ 124 meV. Material parameters given in Ref. [10]: $\mu=0.36\pm m_{e}$, $r_{0}=6.4\pm 0.3$ nm are in a wide interval of encertainty while $\kappa=4.4$ is picked up from other works. Choosing the material parameters so that our theory best fit with the experimental data we have $\mu=0.37m_{e}$, $\kappa=4.3$, and $r_{0}=6.1$ nm. We show the experimental data and theoretical exciton energies in Figs. S-1, S-2, and S-3, which demonstrate the high consistency of our theory with the experiments. Experimental data of Liu et al. [8] are not shown in the figures because they are almost identified with the ones of Ref. Stier2018. Numerically, they match with our theory. Indeed, for monolayer WSe2 , $E_{1s}=-172$ meV, $E_{2s}=-41$ meV, $E_{3s}=-20$ meV, $E_{4s}=-11$ meV. They lead to the transition energies $2s-1s$ 131 meV, $3s-2s$ 21 meV, $4s-2s$ 30 meV. Our theory give -168.38, -38.34, -16.32, and - 8.97 meV for energies, and $2s-1s$ 130.04 meV, $3s-2s$ 22.02 meV, $4s-2s$ 29.37 meV for transition energies.
Figure S-1: Exciton energies for the $ns$ states in monolayers WSe2, WS2, MoSe2, MoS2, and MoTe2 calculated by formula (S-115) (horizontal axis) compared with the experimental data [10, 9] (vertical axis). Material parameters used for calculations are provided in Table S-I.
Figure S-2: Exciton energies for the $ns$ states in monolayers WSe2, WS2, MoSe2, and MoS2 calculated by formula (S-115) (black line) compared with the experimental data [10, 9] (red circle). Material parameters used for calculations are provided in Table S-I.
Figure S-3: Exciton energies for the $ns$ states in monolayer MoTe2 calculated by formula (S-115) (black line) compared with the experimental data [10] (red circle). Material parameters used for calculations are from the best fit.